### v1

# zkSTAR: A zero knowledge system for time series attack detection enforcing regulatory compliance in critical infrastructure networks

Abstract
Industrial control systems (ICS) form the operational backbone of critical infrastructure networks (CIN) such as power grids, water supply systems, and gas pipelines. As cyber threats to these systems escalate, regulatory agencies are imposing stricter compliance requirements to ensure system-wide security and reliability. A central challenge, however, is enabling regulators to verify the effectiveness of detection mechanisms without requiring utilities to disclose sensitive operational data. In this paper, we introduce zkSTAR, a cyberattack detection framework that leverages zk-SNARKs to reconcile these requirements and enable provable detection guarantees while preserving data confidentiality. Our approach builds on established residual-based statistical hypothesis testing methods applied to state-space detection models. Specifically, we design a two-pronged zk-SNARK architecture that enforces (i) temporal consistency of the state-space dynamics and (ii) statistical consistency of the detection tests—allowing regulators to temporally verify alarm correctness without visibility into utility-level data. We formally analyze the soundness and zero-knowledge properties of our framework and validate its practical feasibility through computational experiments on real-world ICS datasets. As a result, our work demonstrates a scalable, privacy-preserving alternative for regulatory compliance for ICS driven critical infrastructure networks.
I Introduction
Critical Infrastructure Networks (CINs)—such as gas, water, and power systems—are large-scale cyber-physical systems managed by multiple stakeholder entities. The growing integration of information and operational technologies (IT-OT) driven by increased IoT adoption has introduced new vulnerabilities to data-driven cyberattacks. Such attacks often target the Industrial Control Systems (ICS) of utility stakeholders, aiming to degrade asset performance and potentially trigger cascading failures across the entire network. Although attack detection is performed locally at stakeholder ICSs, regulatory entities such as Information Sharing and Analysis Centers (ISACs) aggregate alarms and statistics from all utilities in real time to enhance situational awareness, reduce false positives, and expedite decision-making. The effectiveness of ISACs in fulfilling their objectives fundamentally depends on the quality of local ICS detection metrics reported by individual utilities. Inconsistent or low-quality local insights can undermine the ISAC’s ability to construct a reliable network-wide situational awareness model, potentially causing unnecessary service disruptions. Consequently, there remains a critical need for methods that allow regulators to publicly verify local ICS detection outcomes without resorting to exhaustive utility data audits, which are costly, impractical for real-time operations, and intrusive to utility privacy. Therefore, this paper introduces zkSTAR, a framework that enables publicly verifiable detection of data-driven ICS attacks while preserving complete zero knowledge of local ICS state-space dynamics.
The ICS serves as the primary OT backbone for any CIN utility capable of providing control, actuation and data storage capabilities of infrastructure assets in a scalable manner. Data-driven ICS attack detection methods primarily utilize models that characterize the state space of the local utility network which can help unveil deviations from expected steady state conditions using the observed sensor measurements [1]. These deviations or residuals can then be used as inputs for statistical hypothesis tests to distinguish between normal and abnormal operation profiles subject to Type-I and Type-II error parameters of the test [2]. An alarm signifying an attack event is triggered locally when the test reject its null hypothesis indicating statistical inconsistencies between the observed and the expected sensor measurements subject to state space dynamics.
With existing state of the art methods, the only way to conduct regulatory verification of detection outcomes of utility-based state space models is through a full-fledged data audit of ICS operations. For such an audit, the utility would have to relinquish the ownership of their local IoT datasets to a regulator raising significant privacy and data governance hurdles. Next, utility datasets must be shown to be temporally consistent across the time domain. In other words, the regulator must be able to exhaustively certify the state space estimates of successive time steps have been generated from previous estimates. Finally, the regulator must also be able to demonstrate statistical consistency which implies that the correct statistical testing procedure was followed to raise alarms. As a result, regulatory data audits can become extremely expensive and would come at the cost of significant privacy issues for utilities.
Therefore, in this paper, we develop zkSTAR, a framework that is geared towards guaranteeing temporal and statistical consistency for state space driven statistical detection of data driven ICS attacks. At its core, zkSTAR is specifically designed to demonstrate temporal and statistical consistency of ICS systems with highly non-linear state space dynamics in complete zero knowledge. zkSTAR is based on the zkSNARK (zero-knowledge succinct, non-interactive argument of knowledge) paradigm and provides the ability to publicly verify attack detection outcomes for a state space modeling paradigm in a trustworthy fashion. We present a detailed overview of our contributions as follows:
  • We develop a zkSNARK based prediction framework that leverages an extended Kalman filter based model to detect data driven attacks on ICSs of utility stakeholders in complete zero-knowledge of local sensor or state space dynamics.
  • We establish comprehensive theoretical guarantees pertaining to the necessary and sufficient conditions for asserting temporally and statistically consistent regulatory compliance in the context of time-series anomaly detection using non-linear state space models.
  • We design a zkSNARK based kernel decomposition scheme that provides the ability to prove temporal and statistical consistency claims in a computationally efficient manner.
  • We provide a detailed container-based computational testbed for evaluating the performance of zkSTAR with respect to a detailed case studies involving real-world datasets.
II Related Works
Attacks on ICSs that are solely targeted towards the IT systems can be detected and isolated using network traffic monitoring paradigms [3, 4, 5]. However, data-driven attacks are harder to detect through such mechanisms especially in the ICS context owing to their ability to impact critical pathways in both the IT and OT layers [6] potentially causing critical infrastructure damage. Such attacks typically involve manipulation of sensor data so as to effect cascading failures across networked ICS in addition to degrading the performance of critical infrastructure assets. Despite robust network traffic monitoring frameworks, such attacks can go undetected since they involve manipulation at the OT layer[7, 8].
Therefore, several model-based mechanisms have been proposed to detect ICS data-driven attacks [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21]. Among these, the Kalman filter based mechanism of detecting attack proves to be a robust and flexible option [2, 1, 22, 23, 24] owing to its reliance on residuals emanating from a well-formed state space model. Such techniques utilize the residuals obtained from Kalman model predictions and the observed sensor measurements as fundamental inputs to statistical hypothesis tests in order to flag alarms. As a result, a Kalman filter based state space estimation mechanism can help detect anomalies emanating from data driven attacks distinguish between routine faults and attacks using degradation models [22] and use decentralized methodologies to raise network-wide alarms [23].
From the regulatory perspective however, it is essential to not only receive timely notifications regarding attack detection at utility stakeholders but also ensure the integrity of attack detection so as to limit its network-wide impacts and reduce false alarm rates. Given the trepidation of utilities in being able to share data with regulatory agencies due to privacy and logistical hurdles, enabling regulatory compliance of local detection frameworks acquires an entirely new degree of complexity. As a result, it becomes important to ensure that the right sequence of steps has been followed to determine these insights without belying the underlying data. Zero-knowledge proofs form a viable alternative to solve this conundrum with the added capability of complementing privacy mechanisms as well [25].
Zero-knowledge proofs were originally proposed by Goldwasser et al. [26] in which a Prover attempts to convince a Verifier that some mathematical statement is true. The basic premise was that the Verifier must not glean any other information about the secrets held by the Prover that validate a set of mathematical statements. Original propositions of zero-knowledge proofs were interactive, i.e. the Verifier and the Prover exchange a series of messages in order to establish the validity of the mathematical truths. However, Blum et al. [27] showed that zero-knowledge proofs can also be made non-interactive by using a common short random string (CRS). These initial works were followed by establishing the concepts of succinctness that enables efficient verification with minimal overhead [28] which were combined with non-interactivity [29].
In the past decade, there have been several breakthroughs that have paved the way for practical, computationally feasible implementations of zero-knowledge Succinct, Non-Interactive Arguments of Knowledge (zk-SNARKs) [30, 31, 32, 33, 34]. Within the domain of machine learning, zero-knowledge proofs have been used to establish the integrity of training [35, 36, 37] and prove model-ownership [38]. However, there exists a critical gap pertaining to the development of zkSNARKs for time-series anomaly detection using state space dynamics of utility ICS. As a result, considering state of the art methods, providing trustworthy data driven attack detection for fulfilling regulatory compliance requirements in complete zero-knowledge in currently infeasible. In this paper, we develop zkSTAR, a framework that can generate computationally efficient zkSNARKs of data driven attack detection frameworks that are based on Kalman filter driven utility ICS state space models. zkSTAR is primarily geared towards ensuring regulatory compliance in data driven attack detection with full privacy preservation of utility stakeholder data. In fact, zkSTAR is capable of enabling regulatory agencies to verify utility level attack detection mechanisms in zero knowledge of their local state space model altogether.
III Nonlinear State Space Modeling
For developing zkSTAR, we leverage the extended Kalman Filter state space models owing to their flexibility in handling highly non-linear spatial and temporal interdependencies among several ICS components [39]. Therefore, we characterize the EKF-based, utility-level attack detection model as a foundational component of zkSTAR for achieving regulatory compliance under zero knowledge. Within zkSTAR, we develop a state-space formulation capable of identifying attack-induced anomalies through locally executed statistical hypothesis tests.
III-A Extended Kalman Filter Model
We present a brief overview of the Extended Kalman Filter model in this section while a detailed description of the model has been provided in Appendix -A. Let g,h{g},{h} denote the state transition and observation functions respectively, $x_{t}\in R^{m}$ represents the latent space embedding, $u_{t}\in R^{m}$ represents the control action and $y_{t}\in R^{d}$ represents noisy sensor measurements from asset sensors. Additionally, let process and measurement noises follow multivariate normal distributions $N(0,Q_{t})$ and $N(0,R_{t})$ respectively. Now, we define our EKF model based on the following equations:
$\hat{x}_{t|t-1}=g(\hat{x}_{t-1|t-1},u_{t-1}),$ (1)
$r_{t}=y_{t}-h(\hat{x}_{t|t-1}),$ (2)
$P_{t|t-1}=G_{t}P_{t|t-1}G^{T}_{t}+Q_{t-1},$ (3)
$S_{t}=H_{t}P_{t|t-1}H^{T}_{t}+R_{t},$ (4)
$K_{t}=P_{t|t-1}H^{T}_{t}S_{t}$ (5)
$\hat{x}_{t|t}=\hat{x}_{t|t-1}+K_{t}r_{t}$ (6)
$P_{t|t}=(I-K_{t}H_{t})P_{t|t-1}$ (7)
In our EKF model, we have $P_{t|t-1},P_{t|t}$ represent the predicted and the updated covariance estimates respectively, while $S_{t}$ represents the residual covariance at $t$. The state transition and the observation matrices at $t$ given by $G_{t},H_{t}$ respectively are computed using the Jacobians of $g,h$.
III-B Hypothesis Test for Anomaly Detection
We consider $r_{t}$ as our residual and $S_{t}=U_{t}\Sigma_{t}U_{t}^{T}$ as the residual covariance. We denote the standardized vector of PC scores as $\tau_{t}=(S_{t})^{-1/2}(r_{t})=U_{t}\Sigma_{t}^{-1/2}(r_{t})$ such that $\tau_{t}\sim N(0,I)$ [1, 40]. This implies that $\tau_{t}||^{2}_{2}\sim\chi^{2}_{p}$, where $p<d$ is the degrees of freedom corresponding to the number of principal components used [40].
$\rho=\begin{cases}1,&\text{if }T_{\chi^{2},t}=\|\tau_{t}\|^{2}_{2}>\chi^{2}_{m,\alpha},\\ 0,&\text{otherwise.}\end{cases}$ (8)
A generic hypothesis testing procedure applicable to sensor and residual monitoring has been discussed in greater detail in Appendix -B.
We define a measure of probabilistic similarity to help establish the distinctness of two sensor measurement vectors as stated in Definition 1.
Definition 1 (Probabilistic Similarity).
[…]
The concept of probabilistic similarity is particularly useful while handling issues for numerical precision issues wherein the sensor data measurement stored and processed in the ICS differs from the original value recorded by the sensors. We now proceed towards theoretical characterization of the zkSNARKs to enforce temporal and statistically consistent detection outcomes.
IV Regulatory Compliance using zkSNARKs
Detecting attacks requires monitoring covariance matrices of the residuals to drive the underlying $\chi^{2}$ hypothesis tests. However, publicly validating alarms by an external entity such as a regulator would require the disclosure of residual vectors and covariance matrices that enable these hypothesis tests. Therefore, a critical regulatory challenge is the verification of the hypothesis test outcomes as reported by the utilities in the absence of their state space residual and covariance estimates. Zero-knowledge proofs form a viable alternative for solving these critical pain points by helping achieve public verification of utility level attack detection outcomes for regulatory purposes.
Therefore, in this paper, we employ zero knowledge proofs to enable regulatory compliance and verification. We concentrate on a utility-oriented subsystem represented by a state-space model as outlined in Section III. Our privacy considerations specifically address the disclosure risks linked to the public revelation of parameters $C$ and residuals $r_{t}$ derived from the Kalman Filter-based state-space estimation, which are crucial for validating the alarm value $\rho_{t}$. As a result, our methodology specifically leverages zk-SNARKs to help assert temporal consistency and statistical consistency for utility-level $\chi^{2}$ hypothesis tests without the need to divulge state space variables.
IV-A Zero Knowledge Preliminaries
We begin our discussion of zk-SNARK based regulatory compliance schemes for cyber attack detection by establishing four fundamental definitions.
Definition 2 (Hash Function).
A hash function denoted by hash can be defined such that it can consume arbitrary length inputs so as to map to fixed length outputs of size $n$
[…]
Definition 2 portrays a hashing function that will be utilized to ensure the integrity of state space estimates. We assume that the function hash preserves integrity, one-wayness, hiding and collision resistance properties [41].
Definition 3 (Pretrained State Space Model).
The non-linear Kalman Filter based state space can be represented by a pretrained model $\mathcal{M}$ parametrized by $\theta$ such that the following conditions hold
$\mathcal{M}(\Delta^{in}_{t};\theta)\mapsto\Delta^{out}_{t}$
$\Delta^{in}_{t}=(\mathcal{H}^{in}_{t},\hat{x}_{t|t},y_{t}),$
$\Delta^{out}_{t}=(\mathcal{H}^{out}_{t},\rho_{t},\hat{x}_{t+1|t+1},r_{t},K_{t},Q_{t},R_{t})$
[…]
Definition 3 helps define a pretrained LSTM based model which encapsulates the non-linear Kalman Filter based state space as defined in Equations (24)-(30). More specifically, the inputs and outputs of $\mathcal{M}$ at time $t$ are represented by $\Delta^{in}_{t}\text{ and }\Delta^{out}_{t}$ respectively.
Definition 4 (zk-SNARK Setup Phase).
Given a standard security parameter $\lambda$, the zk-SNARK setup phase can be defined by function Setup such that
[…]
Definition 4 discusses the setup phase of the zk-SNARK proof mechanism that involves the generation of $pk$ and $vk$ that represent the proving and verification keys respectively. We assume that the setup phase needs to be executed once for each utility. Such assumptions are driven by the fact that a pretrained, high accuracy state-space model $\mathcal{M}$ does not require frequent updates to the weight parameter. However, we note that the setup function needs to be re-evaluated upon each update of the weight parameter $\theta$.
Definition 5 (zk-SNARK Proving Phase).
The zk-SNARK proof generation phase can be defined by function Prove such that
[…]
Definition 5 details the proving phase of the zk-SNARK. The Prove function consumes the proving key $pk$, $\mathcal{M}$ parametrized by $\theta$ as well as the inputs $\Delta^{in}_{t}$ and outputs $\Delta^{out}_{t}$ at time $t$ to generate a proof artifact $\Pi_{t}$. The Prove function needs to be executed for each timestep $t$ at the utility level.
Definition 6 (zk-SNARK Verification Phase).
The verification of zk-SNARK proof artifacts can be represented by the function Verify such that
[…]
Definition 6 presents the Verify function that aims to enable regulators to verify reported proofs $\Pi_{t}$ based on the verification key $vk$ as well as inputs $\Delta^{in}_{t}$ and $\Delta^{out}_{t}$. The Verify function provides output $\Phi_{t}\in\{0,1\}$ which is indicative of successful verification.
Zero knowledge regulatory compliance for attack detection relies on two fundamental operational attributes of each utility. These fundamental operational attributes pertain to temporal consistency and statistical consistency. We argue that these two fundamental operational attributes are necessary and sufficient conditions to enable zero-knowledge regulatory compliance verification. Our argument is based on the ability of the pretrained model $\mathcal{M}$ to recursively ensure that state space estimates are derived from previous predictions while being statistically consistent with expected residual covariance. In the following subsections, we formally state each operational attribute separately as a precursor for guaranteeing integrity of detection outcomes.
IV-B Temporal Consistency
Temporal consistency pertains to the recursive relationship between a posteriori state space estimates at $t$ and $t+1$ respectively. Temporal consistency constraint encoded in the non-linear Kalman filter model ensures that the posteriori estimate $\hat{x}_{t+1|t+1}$ is a derivative of $\hat{x}_{t|t}$. We state this formally as part of Definition 7.
Definition 7 (Temporal Consistency).
The state space system is considered to be temporally consistent if for each consecutive time step pairs $t,t+1$, the following condition holds.
$x^{out}_{t}=x^{in}_{t+1}$
where,
$x^{out}_{t}=\hat{x}_{t|t-1}+K_{t}r_{t},\text{ and, }\hat{x}_{t+1|t}=g(x^{in}_{t+1},u_{t})$
As a consequence of Definition 7 we can make the temporal consistency claim presented in Theorem IV.1.
Theorem IV.1 (zk-STAR Temporal Consistency Conditions).
Given every consecutive time step pairs $t,t+1$ such that $vk_{t}=vk_{t+1}$, $pk_{t}=pk_{t+1}$, temporal consistency can be asserted if and only if the following conditions hold
$\mathcal{H}^{in}_{t+1}=\mathcal{H}^{out}_{t}\text{ , and }\Phi_{t+1}=1$
Proof.
Proof provided in Appendix -C
Theorem IV.1 shows that the integrity of temporal posteriori predictions from an EKF driven state space system can be established using a zkSNARK proof artifact if and only if the verification keys remain invariant; the hash of the inputs to the EKF model at each time step is equal to the hash of the previous EKF output; and the proof artifact itself can be verified using the verification keys. In other words, Theorem IV.1 demonstrates that if temporally inconsistent values for carrying out the EKF predictions, then either the verification fails for at least one time step or the keys used for verifying the proof artifact will differ. Therefore, to establish temporally consistency from a regulatory perspective, it is necessary and sufficient to check for the invariance of keys, alignment of inputs and outputs and the verification status across successive time steps.
IV-C Statistical Consistency
Consider the local attack detection framework characterized by the $\chi^{2}$ hypothesis test presented in Section III-B with a Type-I and Type-II error rates of $\alpha$ and $\beta$ respectively. The $\chi^{2}$ test relies on the residual and state space estimates provided by $\mathcal{M}$ that is parametrized by $\theta$. As a result of Equation (34), the local attack detection framework is expected to yield an alarm value denoted by $\rho_{t}=1$, when an attack is taking place while maintaining $\rho_{t}=0$ under normal circumstances. We utilize these fundamental principles to establish the notion of statistical consistency. For the purposes of statistical consistency, we denote the genuine and the reported outcomes of the $\chi^{2}$ test as $\rho^{real}_{t}\in\{0,1\}$ and $\rho^{rep}_{t}\in\{0,1\}$ respectively.
Definition 8 (Statistical Consistency).
The local attack detection framework is deemed to be statistically consistent if $\rho^{real}_{t}=\rho^{rep}_{t}$.
Statistical consistency ensures that reported alarm values are consistent with the outcomes of the attack detection frameworks that is based on the $\chi^{2}$ hypothesis test. However, from a real-world deployment perspective, the sensor data measurements observed in the field might have slight inaccuracies as they are transmitted across the utility ICS. These inaccuracies could arise due to sensor calibration errors and numerical precision limitations especially on inexpensive sensor hardware which are commonly found in large scale ICS. In such cases, expecting to acquire and store a copy of the sensor data measurement at different levels of the ICS that is numerically identical to the one that was originally recorded by the sensor is unrealistic. However given the inherent robustness of hypothesis tests towards robustness, utilizing a sensor data measurement that is similar but not necessarily identical to the originally recorded value is sufficient for establishing statistical consistency. As a result, we develop Lemma 1 as a means to formally establish the notion of probabilistic similarity of sensor measurements and its connection to statistical consistency of the zkSTAR framework.
Lemma 1.
Probabilistically similar sensor measurement realizations $\tilde{y}_{t},y_{t}$ are necessary and sufficient conditions for observing statistically coherent outcomes $\tilde{\rho}_{t},\rho_{t}$.
Proof.
Proof provided in Appendix -D
Lemma 1 shows that as long as the sensor signal used to carry out EKF predictions is probabilistically similar to the true value of the sensor signal the zkSTAR framework will continue to deliver statistically consistent alarm detection outcomes. Moreover, with the help of Lemma 1, we can also state that a data driven ICS attack orchestration might be able to succeed only when the malicious sensor measurements are probabilistically similar to the originally recorded sensor measurements. Practically, such attacks end up being inconsequential [22] since they are unable to effect any material damage to the ICS assets in the first place. Therefore, Lemma 1 serves as the basis for Theorem IV.2 which derives an integral link between statistical and temporally consistency on one hand and the key invariance and verification status on the other.
Theorem IV.2 (zk-SNARK Statistical Consistency Conditions).
Given consecutive time step pairs $t,t+1$ with $(vk_{t},pk_{t})$ and $(vk_{t+1},pk_{t+1})$ and a temporally consistent system, statistical consistency can be guaranteed if and only if $\Phi_{t+1}=1$
Proof.
Proof provided in Appendix -E
Theorem IV.1 shows that alarm values obtained from zkSTAR can be deemed statistically consistent if they share the same proving and verification keys, resulting in a successful verification while adhering to Theorem IV.1. As a result, a regulatory agency can reject the veracity of reported alarms from utility stakeholder in spite of successful verification if either of the conditions laid down in Theorem IV.2 are not met. Additionally, Theorem IV.2 shows that unsuccessful verification of stakeholder proofs implies a statistical inconsistency in alarm detection despite fulfilling temporal consistency claims. Lastly, Theorem IV.2 also establishes that temporal inconsistencies necessarily indicate a statistically inconsistent system regardless of the alarm and verification status.
V Computational Architecture of zkSTAR
The algorithmic framework of zkSTAR is based on a temporal state space modeling paradigm that is driven by a non-linear LSTM-based Kalman filter model. There are three fundamental challenges that arise when attempting to detect data-driven attacks in ICS in zero knowledge. First, asserting temporal consistency between successive state space estimations is challenging due to the highly non-linear and dynamic nature of utility based ICS subsystems. Second, ensuring the statistical consistency of the detection mechanism without revealing parameters of the statistical test is difficult. Lastly, the nature of high-fidelity streaming sensor datasets possess a fine temporal granularity that leads to an additional layer of complexity when characterizing attacks.
As a result, the zkSTAR framework is fundamentally designed to ensure temporal consistency (TC) and statistical consistency (SC) in the context of high-fidelity streaming sensor data. In order to address the challenge of high fidelity sensor information the zkSTAR framework adopts a temporal decomposition scheme that is specifically designed to cater to TC and SC aspects. Therefore, in this section, we provide an in-depth overview of the temporal decomposition scheme before proceeding with a technical discussion and algorithmic review of TC and SC components of zkSTAR.
V-A Temporal Decomposition Scheme
We consider a real-time attack detection horizon that is discretized into a sequence of detection windows denoted by set $\mathcal{W}_{i},\forall i\in\{1,2,\ldots\}$. Each detection window $\mathcal{W}_{i}$ is a set of $W$ time intervals corresponding to utility ICS state space model. Each time interval is represented by $w_{ij},\forall j\in\{1,W\}$ and is primarily geared towards enforcing temporal consistency across multiple prediction time steps. Every TC interval is denoted by $w_{ij},i\in\{1,2,\ldots\},j\in\{1,W\}$ and consists of $D$ consecutive timestamps such that $w_{ij}=\{t_{w_{ij}},t_{w_{ij}}+1,t_{w_{ij}}+2\ldots t_{w_{ij}}+D\}$, where $t_{w_{ij}}$ is the starting timestamp of detection window $w_{ij}$. The relationship between each detection window $\mathcal{W}_{i}$ and its constituent TC intervals $w_{ij}$ can be compactly denoted using the following equations.
$\mathcal{W}_{i}=\{w_{i1},w_{i2},\ldots w_{iW}\},\ \forall i\in 1,2,\ldots$ (9)
$w_{ij}=\{t_{w_{ij}},t_{w_{ij}}+1,t_{w_{ij}}+2\ldots t_{w_{ij}}+D\},\ \forall j\in\{1,W\}$ (10)
Without loss of generality, let us consider a detection window $\mathcal{W}$ and its constituent TC intervals $w_{j}\in\mathcal{W},\ j\in\{1,W\}$. We know from Section III that for any time step $t_{w}+k\in w$, $r_{t_{w}+k}\sim N(0,S_{t_{w}+k})$ where $r_{t_{w}+k},S_{t_{w}+k}$ are the residual and the covariance matrix predicted by the pre-trained EKF model for time step $t_{w}+k$. Since $r_{t_{w}+k}\sim N(0,S_{t_{w}+k})$, the residuals and covariance matrices can be aggregated across all TC intervals for detection window to yield $r_{W}$ and $S_{W}$ as represented in Equation (11).
$r_{W}=\sum\limits_{j=1}^{W}\sum\limits_{k=1}^{D}r_{t_{w_{j}}+k}\ \text{,}\ S_{W}=\sum\limits_{j=1}^{W}\sum\limits_{k=1}^{D}S_{t_{w_{j}}+k}$ (11)
As a result of hierarchical decomposition, there exists a distinct detection alarm $\rho_{W}$ for each detection window $\mathcal{W}$.
The temporal decomposition scheme paves the way to leverage zkSNARKs that can help enforce integrity of state space predictions based on a pre-committed EKF model. However, the zkSNARK circuit size grows exponentially with the growth in the detection window size making the proof generation step computationally expensive especially for larger systems. Therefore, in order to achieve computational efficiency, while retaining the practicality of zkSNARKs, we adopt a kernel based strategy that decouples compute tasks pertaining to an EKF based state space attack detection framework. Specifically, we design zkSNARK based kernels that target individual sub tasks pertaining to temporal consistency of EKF state space predictions as well as statistically consistent outcomes of the detection model that operates on batched EKF residual outputs.
We also note that state of the art zkSNARK frameworks currently do not support linear algebra operations, such as linear solvers, SVD decompositions while lacking the capability of evaluating derivatives of complex sub-modules. This lack of support largely emanates from the significant computational cost of conducting these operations in full zero knowledge making them impractical. In order to circumvent these constraints, we adopt a compute-offline-validate-online approach wherein computationally expensive steps such as derivative evaluations, linear system solutions and SVD decompositions are carried out off-circuit but are evaluated for consistency with the state space model within the zkSNARK circuit.
Figure 1: Temporal consistency handling in zkSTAR
V-B zkSNARK based Temporal Kernels
For ensuring temporally consistent outcomes, our kernel design strategy involves the breaking down distinct components of the EKF model inference. Specifically, we decompose the EKF inference over a single TC interval into distinct phases that pertain to validating the consistency of the Kalman gain matrix with the state covariance and process and measurement Jacobians; inference of the pre-committed EKF model on one timestamp of the TC interval; and accumulating the residual and its covariance sum across the entire TC interval.
Kalman Reconstruction Check (KRC) Kernel: The EKF model presented in Section III relies on successive linear solves of Equation (28) in order to determine the Kalman gain $K_{t}$. However, this poses a major challenge since implementing linear system solvers within a zkSNARK is computationally very expensive. Therefore, we develop a computational kernel tailored for the zkSNARK paradigm that demonstrates that offline computed matrices Kalman gain $K$, LSTM jacobian $H$, residual covariance $S$ and state covariance $P$ are consistent with respect to the state space EKF model. The result is the KRC kernel depicted in Algorithm 1 which yields a boolean output $\kappa_{t}$ depending on whether the error in Equation (28) is below or above threshold $\epsilon_{KC}$
Algorithm 1 zkSTAR Kernel for Kalman Gain Reconstruction
1:function KRC-Kernel($K$, $S^{-1}$, $H$, $P$, $\epsilon$)
[…]
3:if $\|KS-PH^{\top}\|_{2}^{2}<\epsilon$ then $\kappa\leftarrow 1$
4:end if
5:return $\kappa$
6:end function
Extended Kalman Filter (EKF) Kernel: For enforcing the state space dynamics at each timestamp in zero knowledge, we develop a dedicated zkSNARK based computational kernel for EKF denoted by $\mathcal{M}^{EKF}_{t}$.
[…]
Fundamentally, $\mathcal{M}^{EKF}$ forms a mapping between input tuple $\mathcal{I}^{EKF}_{t}$ and output tuple $\mathcal{O}^{EKF}_{t}$. The input tuple in turn consists of $\mathcal{X}_{t-1}=\{x_{t-1|t-1},P_{t-1|t-1}\}$ which denotes a collection of state and covariance estimates from the prior time step. $\mathcal{J}_{t}=\{G_{t},H_{t},K_{t}\}$ represents a collection of the Jacobian matrices of the LSTM and feed-forward neural layers in the EKF formulation along with the Kalman gain matrix that are computed offline. Lastly, $y_{t}$ represents the sensor data sampled at time stamp $t$. The output tuple $\mathcal{O}^{EKF}_{t}$ consists of $r_{t}$$S_{t}$$\kappa_{t}$, which denote the residual and its covariance as well as the status of the Kalman Gain reconstruction check.
Algorithm 2 zkSTAR Kernel for Extended Kalman Filter
1:function EKF-Kernel($x_{t-1|t-1}$$P_{t-1|t-1}$$y_{t}$$G_{t}$$H_{t}$$K_{t}$)
2: compute $\hat{x}_{t|t-1}$ using Equation (24).
3: compute $r_{t}$ using Equation (25)
4: compute priori covariance $P_{t|t-1}$ using Equation (26)
5: compute inverse innovation $S_{t}^{-1}$ using Equation (27).
[…]
7: update posteriori estimate $\hat{x}_{t|t}$ using Equation (29)
8: update the state posteriori covariance $P_{t|t}$ using (30) return $r_{t}$$S_{t}$$\kappa_{t}$
9:end function
Algorithm 2 represents the computational kernel for EKF within the zkSTAR framework denoted by EKF-Kernel. For a specific timestamp $t$, EKF-Kernel computes the priori state and covariance estimates $\hat{x}_{t|t-1},P_{t|t-1}$ as well as the resulting residuals $r_{t}$ using Equations (24), (26) and (25) respectively. Using Equation (27), the kernel estimates the inverse of the innovation matrix or the residual covariance $S_{t}^{-1}$. Next, the KRC-Kernel is invoked to validate the consistency of the Kalman gain matrix $K_{t}$ with respect to the Jacobians $G_{t}$, $H_{t}$ as well as the posteriori covariance $P_{t|t}$. Subsequently, the posteriori state and covariance estimates are updated according to Equations (29) and (30). The kernel returns the current residual $r_{t}$, the residual covariance $S_{t}$ and the Kalman gain reconstruction status $\kappa_{t}$.
Temporal Consistency (TC) Kernel: While the EKF kernel can enforce state space dynamics for each timestamp $t$, it cannot independently ensure the integrity of state space estimates across multiple time stamps. In order to do so, we need a mechanism to assert the integrity between distinct timestamps. Therefore, we develop a temporal consistency (TC) kernel that is geared towards ensuring the integrity across dictinct timestamps. Using the temporal decomposition scheme highlighted in Section V-A we focus on a single TC interval denoted by $w$ to aid our discussion. We denote the inputs to the TC kernel in terms of three sets: $\mathcal{X}^{TC}_{w-1},\mathcal{R}^{TC}_{w-1}$ and $\mathcal{K}^{TC}_{w-1}$ which represent state, residual and KRC status estimates from the previous TC interval w1w-1.
$\mathcal{X}^{TC}_{w-1}=\underbrace{\{(x_{w-1|w-1}||N^{x}_{w-1}),(P_{w-1|w-1}||N^{P}_{w-1})\}}_{\text{nonce appended state estimates}}$ (13)
$\mathcal{R}^{TC}_{w-1}=\underbrace{(r_{w-1}||N^{r}_{w-1}),(S_{w-1}||N^{S}_{w-1})}_{\text{nonce appended residual and covariance}}$ (14)
$\mathcal{K}^{TC}_{w-1}=\underbrace{(\kappa_{w-1}||N^{\kappa}_{w-1})}_{\text{nonce appended KRC status}}$ (15)
Each set in turn consists of the raw values that are appended with a unique randomly chosen none value to provide a good degree of entropy to the resulting hash strings. As a result, $\mathcal{X}^{TC}_{w-1}$ consists of $(x_{w-1|w-1}||N^{x}_{w-1}),(P_{w-1|w-1}||N^{P}_{w-1})$ where $x_{w-1|w-1}$, $P_{w-1|w-1}$, $N^{x}_{w-1}$ and $N^{P}_{w-1}$ denote the posteriori state and covariance estimate and their corresponding nonce values from the previous TC interval. Similarly, $\mathcal{R}^{TC}{w-1}$ comprises the residual and covariance pairs $(r{w-1}||N^{r}_{w-1})$ and $(S_{w-1}||N^{S}_{w-1})$, where $r_{w-1}$ and $S_{w-1}$ represent the innovation and its associated covariance from the previous window, and $N^{r}_{w-1}$ and $N^{S}_{w-1}$ are the corresponding nonce values ensuring stochastic uniqueness of their hash representations. Finally, $\mathcal{K}^{TC}{w-1}$ encapsulates $(\kappa_{w-1}||N^{\kappa}_{w-1})$, where $\kappa{w-1}$ denotes the binary KRC indicator and $N^{\kappa}_{w-1}$ is its appended nonce value used to prevent deterministic linkage across temporal proofs.
$\mathcal{Z}^{TC}_{t}=\{\underbrace{(y_{t}||N^{y}_{t}),(G_{t}||N^{G}_{t}),(H_{t}||N^{H}_{t}),(K_{t}||N^{K}_{t})}_{\text{nonce appended sensor data, Jacobian and Kalman Gain}}\}$ (16)
$\mathcal{Z}^{TC}_{w}=\{{\underbrace{{\mathcal{Z}^{TC}_{t_{w}},\mathcal{Z}^{TC}_{t_{w}+1},\ldots\mathcal{Z}^{TC}_{t_{w}+D}}}_{\text{EKF inputs for TC interval w}}}\}$ (17)
Additionally, the TC kernel also utilizes the set $\mathcal{Z}^{TC}_{w}$ which encapsulates the offline computed inputs and sensor data denoted by $\mathcal{Z}^{TC}_{t}$ meant for the EKF Kernel invocations at each timestep of the TC interval. The set $\mathcal{Z}^{TC}_{t}$ contains the nonce-appended sensor measurements and model parameters $(y_{t}||N^{y}_{t})$, $(G_{t}||N^{G}_{t})$, $(H_{t}||N^{H}_{t})$, and $(K_{t}||N^{K}_{t})$ corresponding to the sensor data, the Jacobians, and Kalman gain, respectively. Each element in $\mathcal{Z}^{TC}{t}$ is locally randomized through its associated nonce to ensure that the resulting hashed representations remain unlinkable across successive EKF invocations. Collectively, the windowed set $\mathcal{Z}^{TC}_{w}$ aggregates these per-timestep inputs over the duration of the TC interval $w$, thereby serving as the complete batch of EKF inputs required for evaluating the zkSTAR TC kernel and generating the proof of temporal consistency for that interval. Consequently, Equation (18) is used to represent the TC kernel as a mapping denoted by $\mathcal{M}^{TC}$.
$\mathcal{M}^{TC}:\;\underbrace{\mathcal{I}^{TC}_{t_{w}}=\{[\mathcal{X},\mathcal{R},\mathcal{K}]_{w-1},[\mathcal{Z},\mathcal{N}]_{w}\}}_{\text{\small input for TC interval $w$}}\;\mapsto$ (18)
$\qquad\underbrace{\mathcal{O}^{TC}_{t_{w}}=\{[\mathcal{X},\mathcal{R},\mathcal{K}]_{w}\}}_{\text{\small output of TC interval $w$}}.$
Algorithm 3 zkSTAR Temporal Consistency (TC) Kernel
1:function TC-Kernel($\mathcal{X}^{TC}_{w-1}$$\mathcal{R}^{TC}_{w-1}$$\mathcal{Z}^{TC}_{w}$$\mathcal{N}^{TC}_{w}$)
2:#Strip nonce vectors from inputs
[…]
6:for $t\in\{t_{w},t_{w}+1,t_{w}+2\ldots t_{w}+D\}$ do
[…]
8: accumulate residual sum $r_{w}\leftarrow r_{w}+r_{t}$
9: accumulate residual covariance sum $S_{w}\leftarrow S_{w}+S_{t}$
10: accumulate KRC status $\kappa_{w}\leftarrow\kappa_{w}\circ\kappa_{t}$
11:end for
12:#Append respective nonce strings from $\mathcal{N}^{TC}_{w}$
[…]
16:return $\mathcal{X}^{TC}_{w},\mathcal{R}^{TC}_{w},\mathcal{K}^{TC}_{w}$
17:end function
Algorithmically, the kernel is represented using the TC-Kernel function presented in in Algorithm 3. The kernel consumes the outputs of the previous TC interval w1w-1 and iterates through the timestamps $\{t_{w}\ldots t_{w}+D\}$ belonging to the current TC interval $w$. For each timestamp, the kernel strips the nonce vectors to obtain the underlying values which are fed to EKF-Kernel. The residual, covariance and KRC status outputs of the EKF kernel are accumulated across all the timestamps of the current TC interval $w$ through summation and multiplication operations respectively. At each iteration of the kernel, Finally, the kernel returns as the overall outputs of the TC kernel itself.
Figure 2: Statistical consistency estimation in zkSTAR
V-C zkSNARK based Statistical Kernels
Applying the $\chi^{2}$ test on the accumulated residuals and covariance requires us to carry out SVD decomposition to evaluate the test statistic and the alarm criteria. Conducting the SVD decomposition, evaluating the corresponding PC scores, evaluating the test statistic and the subsequent alarm value must be conducted completely in zero knowledge. Therefore, we design a sequence of statistical zkSNARK kernels to help achieve efficiency in zkSTAR.
SVD Reconstruction Kernel: We develop a kernel to validate the SVD decomposition of the accumulated covariance matrix in zero-knowledge. The SVD step is essential to compute the normalized PC scores so as to determine the value of the test statistic in conjunction with the accumulated residual values for each detection window. The SVD validation kernel SVD-Kernel is presented in Algorithm 4.
Algorithm 4 zkSTAR Kernel for SVD Reconstruction
1:function SVD-Kernel($U$, $\Sigma^{-\frac{1}{2}}$,$S$, $\epsilon$)
[…]
3:if $\|(U\Sigma^{-\frac{1}{2}})^{T}(U\Sigma^{-\frac{1}{2}})-S^{T}S\|_{2}^{2}<\epsilon$ then $\eta\leftarrow 1$
4:end if
5:return $\eta$
6:end function
Specifically the SVD kernel consumes the offline computed value of the left singular matrix $U$ as well as the square root of the singular values $\Sigma^{-\frac{1}{2}}$, an error tolerance threshold $\epsilon$ in addition to the accumulated covariance matrix $S$ itself. The kernel returns the validation status $\eta$ depending on whether the SVD reconstruction represented by $U(\Sigma^{-\frac{1}{2}})^{T}(U\Sigma^{-\frac{1}{2}})$ falls within an $\epsilon$ ball of the predicted covariance matrix accumulated across the detection window.
Hypothesis Test Kernel: We conduct the $\chi^{2}$ hypothesis test using the SC-Kernel which considers the accumulated residual and covariance values at the end of every detection window. Additionally, we compute the SVD of the accumulated covariance matrix in an offline manner and validate its consistency using the SVD reconstruction kernel denoted by SVD-Kernel. More specifically, we represent the inputs to the SVD kernel in terms of $\mathcal{Z}^{SVD}_{W}$ as stated in Equation (19) which consists of the singular vectors and the square roots of the singular values.
$\mathcal{Z}^{SVD}_{W}=\{\underbrace{(U_{W}||N^{U}_{W}),(\Sigma^{-\frac{1}{2}}_{W}||N^{\Sigma}_{W})}_{\text{nonce appended SVD values}}\}$ (19)
The hypothesis testing kernel pertaining to detection window W can therefore be denoted as a mapping $\mathcal{M}^{SC}_{W}$ as stated in Equation (20) which maps input set $\mathcal{I}^{SC}_{W}$ to output $\mathcal{O}^{SC}_{t_{w}}$.
$\mathcal{M}^{SC}:\underbrace{\mathcal{I}^{SC}_{W}=\{\mathcal{R}_{W},\mathcal{Z}^{SVD}_{W}\}}_{\text{input for HT}}\mapsto\underbrace{\mathcal{O}^{SC}_{t_{w}}=\{\rho_{W},\eta_{W},\kappa_{W}\}}_{\text{output of HT}}$ (20)
In Equation (20), $\mathcal{R}_{W}$ represents outputs of TC-Kernel obtained upon the culmination of the current detection window while $\mathcal{Z}^{SVD}_{W}$ represents the offline computed SVD of the accumulated residual covariance matrix.
Algorithm 5 zkSTAR Kernel for the $\chi^{2}$ Hypothesis Test
1:function SC-Kernel($\mathcal{R}_{W}$$\mathcal{Z}^{SVD}_{W}$)
2:#Strip nonce vectors from inputs
[…]
5:#Set UCL and default alarm values for W
[…]
9:if $T_{\chi^{2},W}>\chi^{2}_{m,\alpha}$ then $\rho_{W}\leftarrow 1$ then
10:end if
11:return $\rho_{W}$, $\eta_{W}$, $\kappa_{W}$
12:end function
In SC-Kernel represented in Algorithm 5, we consume $\mathcal{R}_{W}$ and $\mathcal{Z}^{SVD}_{W}$ and initiate the hypothesis testing procedure by stripping the nonce vectors from the inputs. We compute the test statistic $T_{\chi^{2},W}$ by utilizing the offline computed SVD values. Additionally, the value of the upper control limit $T^{UCL}_{W}$ is baked into the proving circuit for individual subsystems since it is assumed to remain invariant across time, while the alarm value $\rho_{W}$ is set to 0 by default. We compare the test statistic with $T^{UCL}$ to determine the value of $\rho_{W}$ within the kernel. Lastly, the SVD kernel is invoked to ensure the consistency of the offline computed SVD values with the accumulated sensor covariance matrix $S_{W}$ with a threshold of $\epsilon_{SVD}$ to yield a consistency status denoted by $\eta_{W}$. The kernel outputs a tuple of values consisting of the alarm $\rho_{W}$, the SVD status check $\eta_{W}$ as well as the cumulative KCR status check of $\kappa_{W}$.
V-D zkSTAR Algorithmic Framework
We describe the overall zkSTAR algorithmic framework in Algorithm 6 that encapsulates the temporal and statistical consistency kernels. The framework primarily executes on a rolling horizon of detection windows characterized by $\mathcal{W}_{i}$ with each window consisting of $W$ TC intervals of $D$ timestamps each.
As a result, zkSTAR produces $W$ TC and 1 SC proof artifact for each detection window $\mathcal{W_{i}}$ which is denoted by the set $\mathcal{P_{W_{i}}}$ described in Equation (21).
$\mathcal{P_{W_{i}}}=\Big[\Big(\Pi^{TC}_{w_{i1}}\ldots\Pi^{TC}_{w_{iW}}\Big),\Pi^{SC}_{\mathcal{W_{i}}}\Big]$ (21)
[…]
Algorithm 6 zkSTAR Algorithmic Framework
1:#Set visibility for TC circuit
[…]
3:
4:#Setup proving circuit for TC
[…]
6:
7:#Set visibility for SC circuit
[…]
9:
10:#Setup proving circuit for SC
[…]
12:
13:distribute $vk^{TC},vk^{SC}$ publicly
14:
15:for $i=1,2\ldots$ do $\triangleright$ detection window $\mathcal{W}_{i}$
16:for $j=1,2,\ldots W$ do $\triangleright$ TC interval $w_{ij}$
17: generate nonce values $\mathcal{N}^{TC}_{j}$
18: compute $\mathcal{Z}^{TC}_{j}$ using Equation (17)
19: compute $\mathcal{X},\mathcal{R},\mathcal{K}]_{j-1}$ using Equations (13)-(15).
[…]
22:
23:#Generate TC proof for interval $j$
[…]
25:
26: extract $\mathcal{X},\mathcal{R},\mathcal{K}]_{j}$ from $\mathcal{O}^{TC}_{j}$
27:end for
28: extract $S_{W}$ from $\mathcal{R}_{W}$
[…]
30: generate nonce values $\mathcal{N}^{SVD}$
31: compute $\mathcal{Z}^{SVD}_{W}$ using SVD values $U_{W},\Sigma_{W}^{-\frac{1}{2}},\mathcal{N}^{SVD}$
[…]
33:
34:#Generate SC proof for window $\mathcal{W}_{i}$
[…]
36:
37:return $\mathcal{P_{W_{i}}}=\Big[\Big(\Pi^{TC}_{w_{i1}}\ldots\Pi^{TC}_{w_{iW}}\Big),\Pi^{SC}_{\mathcal{W_{i}}}\Big]$
38:end for
For each detection window $\mathcal{W}i$, the algorithm iterates through the TC intervals wijw{ij} of length $W$. Within each TC interval, a fresh set of nonce values $\mathcal{N}^{TC}{j}$ is generated to enable efficient hiding for hash outputs. Using these nonces, the algorithm constructs the nonce-appended input set $\mathcal{Z}^{TC}{j}$ according to (17), which encapsulates sensor data, model Jacobians, and Kalman gain parameters. The posterior state, residual, and kernel reconstruction variables from the previous interval, $\mathcal{X},\mathcal{R},\mathcal{K}]{j-1}$, are then obtained from (13)–(15). Together, these form the TC-kernel input $\mathcal{I}^{TC}{j}={[\mathcal{X},\mathcal{R},\mathcal{K}]{j-1},[\mathcal{Z}^{TC},\mathcal{N}^{TC}]{j}}$, which is passed to the TC-Kernel for zkSNARK-based verification of temporal consistency. The kernel outputs the updated tuple $\mathcal{O}^{TC}{j}={[\mathcal{X},\mathcal{R},\mathcal{K}]{j}}$, whose elements are extracted and propagated to the next interval.
After all $W$ TC intervals within $\mathcal{W}_{i}$ are processed, the accumulated residual set $\mathcal{R}_{W}$ is used to derive the composite covariance $S_{W}$. An offline singular-value decomposition is then performed to obtain $U_{W}$ and $\Sigma_{W}^{-\frac{1}{2}}$, which serve as normalized orthogonal bases for statistical calibration. A new nonce sequence $\mathcal{N}^{SVD}$ is generated, and the nonce-augmented SVD feature set $\mathcal{Z}^{SVD}_{W}$ is constructed accordingly. Finally, the SC-Kernel consumes $\mathcal{R}_{W}$ and $\mathcal{Z}^{SVD}_{W}$ to compute the high-level decision metrics $\rho,\eta,\kappa]_{W}$, representing the detection confidence, threshold statistics, and overall temporal-consistency verdict for the window $\mathcal{W}_{i}$. This output constitutes the verifiable zkSNARK-backed anomaly indicator used by zkSTAR for secure, decentralized model validation.
VI Experimental Results
Figure 3: HAI Dataset: Detection Quality with varying D and fixed PSF
Figure 4: ORNL-PS Dataset: Detection Quality with varying D and fixed PSF
Figure 5: HAI Dataset: Detection Quality with varying PSF and fixed D
Figure 6: ORNL-PS Dataset: Detection Quality with varying PSF and fixed D
For experimental studies, we utilize the HAI [42] and ORNL-PS [20, 43] dataset. To build a asset level anomaly detector, we generate pretrained non-linear models using the EKF approach to accurately reflect the state-space of the critical asset subsystems as described in Section III. We encode the temporal and statistical components of these models in corresponding zk-SNARKs as described in previous sections. As a result, the zkSTAR framework yields that encode temporal and statistical consistency with respect to detection outcomes.
We carry out detection based on a pre-determined value of upper control limit (UCL) that is representative of the levels of significance used for the $\chi 2$ hypothesis test. These UCL values are generally dependent on the subsystem data and might also rely on engineering domain expertise as well. We use a value of 5e35e-3 and 2e22e-2 for HAI and ORNL-PS datasets that were obtained after rigorous empirical studies.
System Implementation Details: All experiments were carried out on a virtual machine (VM) running Ubuntu 24.04 with 100GB of RAM and 16 vCPUs using Python 3.11 with the detection model inference using PyTorch 2.7.1. We utilize the ezkl library for generating all the zkSNARKs associated with the zkSTAR frameworkWe provide a native as well as a container based mechanism for executing the zkSTAR framework. The native method is primarily meant for comprehensive analysis on a high-performance computing (HPC) clusters to evaluate detection accuracy and computational correctness with varying value of residual aggregation window (referred to as the D value) and precision scale factor (PSF value). The container based method is primarily meant to evaluate real-world system performance in terms of CPU utilization and memory consumption for the distinct processes of witness generation as well as temporal consistency and statistical consistency proofs. We utilize separate containers to represent utility as the prover and regulator as the verifier with an interactive REST API driven interface for evaluation and deployment.
VI-A Detection Quality of zkSNARK outcomes
For assessing the detection quality, we carry out two sets of experiments that depict the effects of varying the PSF with a fixed D and vice versa as well. Specifically, these experiments pertain to a total detection run-time window of 10 minutes in which the attack starts roughly in the middle. Additionally, the attack duration window includes the actual attack occurrence as well as a buffer time of 3 minutes for the HAI dataset. For the ORNL-PS dataset, the attack duration window lasts until the end of the detection run-time window of 10 minutes. Even though the HAI attack scenario lasts 164 secs, we include a buffer time for HAI due to the highly non-linear nature of the detection model that can induce slight delays before anomalies can acquire statistical significance. For these experiments, we consider PSF values of 8,10 and 12, while using 4, 8, 16 and 32 as the D values.
Figures 3 and 4 demonstrate the detection quality with varying D and while fixing the PSF values to 8, 10 and 12 respectively. From both figures we can observe that the detection quality with respect to the UCL remains robust across varying values of D and PSF. Generally, we can observe from both figures that the curves for different D values become more clustered around the non-zkSNARK version with increasing PSF values. This trend remains consistent across both datasets as well. In each case we can observe that the detection consistently occurs within the designated attack duration window providing similar quality as the non-ZKSNARK version.
In Figures 5 and 6, we depict the detection quality with respect to varying PSF values for distinct values of D. In this case we can observe that the lower PSF values usually exhibit greater volatility in detection, while higher PSF values correspond to a less volatile nature of the test statistic. Additionally, we also observe a slight increase in volatility with increasing D which is expected since the errors in residual computation compound for lower PSF values with higher D values.
From these experiments, we can conclude that a higher PSF value might be more desirable for a lower D. However, a mid-range PSF value might also suffice for a higher D value. Regardless, the experiments establish the robustness in detection quality of the zkSTAR framework. The setup times for HAI and ORNL-PS dataset has been provided in Tables III and IV respectively.
Figure 7: Computational error check for SVD of Kalman gain and aggregated covariance matrices
Figure 8: Computational error check for SVD of Kalman gain and aggregated covariance matrices
Dataset PK VK PF
HAI 16011.879 (2271.245) 3.878 (1.014) 0.039 (0.004)
ORNL-PS 16516.104 (2735.173) 4.103 (1.221) 0.040 (0.006)
Table I: Proof (PF), Proving Key (PK), and Verification Key (VK) sizes for Temporal Consistency (mean(std) in MB)
Dataset PK VK witness Circuit Size PF
HAI 12931 3.002 0.023 0.009 0.037
ORNL-PS 12931 3.002 0.016 0.008 0.035
Table II: Proof (PF), Proving Key (PK), and Verification Key (VK) sizes for Statistical consistency (mean(std) in MB)
Setup Setup Time (s)
Phase D=1 D=4 D=8 D=16
Model Export 24.77 3.25 7.96 24.21
Settings Generation 30.02 3.53 9.55 28.99
Circuit Compilation 24.46 2.24 7.06 24.90
Circuit Setup 171.63 146.84 162.04 167.55
Table III: HAI Setup Time Breakdown (in seconds)
Setup Setup Time (s)
Phase D=1 D=4 D=8 D=16
Model Export 24.77 3.25 7.96 24.21
Settings Generation 30.02 3.53 9.55 28.99
Circuit Compilation 24.46 2.24 7.06 24.90
Circuit Setup 171.63 146.84 162.04 167.55
Table IV: ORNL-PS Setup Time Breakdown (in seconds)
D HAI ORNL-PS
Value Witness Circuit Witness Circuit
1 0.197(0.001) 2.94 0.176(0.002) 2.93
4 0.073(0.001) 0.78 0.064(0.001) 0.78
8 0.114(0.001) 1.50 0.101(0.001) 1.49
16 0.197(0.001) 2.94 0.176(0.002) 2.93
32 0.362(0.002) 5.82 0.325(0.004) 5.80
Table V: Temporal Consistency Witness Size with Varying D (mean(std) in MB)
PSF Hypothesis Test D=1 D=4 D=8 D=16
Value Time (s) Time (s) Speedup Time (s) Speedup Time (s) Speedup Time (s) Speedup
8 56.06(22.60) 90.83(5.41) 1.00 80.81(4.19) 4.50 70.61(2.87) 10.29 78.25(3.75 18.57
10 54.26(19.17) 76.16(1.55) 1.00 71.32(3.22) 4.27 75.97(8.88) 8.02 77.11(5.01) 15.80
12 53.49(6.60) 76.74(3.05) 1.00 66.50(2.34) 4.62 74.59(3.19) 8.23 77.79(3.90) 15.78
Table VI: HAI Dataset: Speedup and Proof Generation Times (mean (std) in seconds)
PSF Hypothesis Test D=1 D=4 D=8 D=16
Value Time (s) Time (s) Speedup Time (s) Speedup Time (s) Speedup Time (s) Speedup
8 49.01 (9.68) 73.40 (1.92) 1.00 57.75 (1.35) 5.08 67.69 (1.74) 8.67 72.64 (1.77) 16.17
10 50.79 (1.32) 74.00 (2.06) 1.00 65.06 (1.17) 4.55 68.66 (1.75) 8.62 73.28 (1.83) 16.16
12 52.40 (37.36) 99.94 (1.87) 1.00 64.45 (1.26) 6.20 67.21 (1.22) 11.90 73.32 (2.17) 21.81
Table VII: ORNL-PS Dataset: Speedup and Proof Generation Times (mean (std) in seconds)
VI-B Computational Correctness of zkSNARK outputs
The SVD and Kalman gain decompositions in the zkSTAR framework are computed offline. As a result, the zkSTAR framework outputs the reconstruction error, the associated threshold and a compute check flag for the SVD and Kalman gain matrices. We define the computational correctness as the absolute error emanating from the in-circuit reconstruction of the SVD and Kalman gain matrices computed with respect to the circuit encoded thresholds. Therefore, these experiments demonstrate the variation in relative absolute errors of all matrix elements with respect to the given threshold. In our experiments the threshold values for Kalman and SVD checks were set to 0.1.
Figure 7 denotes box plot series of relative error for varying D values for HAI and ORNL-PS datasets with PSF=12 pertaining to the SVD of the predicted covariance matrices. We note that in case of both datasets, the relative error remains well below 1 for a majority of the cases. This indicates that the reconstruction error remains less than 0.1 in most cases. We observe higher incidence of relative error values going above 1 during the attack window which can be due to the greater volatility in residual values computed during an attack. However, even in the worst case, the relative error with respect to matrix elements does not exceed 1.4 and 1.2 respectively for HAI and ORNL datasets.
Figure 8 denotes scatter plot series of relative errors for varying D values. We note that unlike the SVD check, the Kalman gain check is carried out after each D time steps, resulting in a staggered trend for the first few iterations. The Kalman gain check for both datasets conclusively demonstrates a consistently stable and low error value in the neighborhood of 1e51e-5 for HAI dataset. However, for the ORNL-PS dataset, we see a higher but highly stable relative error values aroun 1e21e-2.
VI-C System Performance of Utility Prover
We begin by analyzing the system performance of zkSTAR proving system. Specifically, we examine the witness, proof, key and circuit sizes in addition to analyzing the utility’s system performance of witness proof generation.
Analysis of Proof and Key Sizes: Tables I and II presents the sizes of the proof, proving and verification keys for the temporal consistency and statistical consistency proofs. Table II also presents the witness size pertaining to the statistical consistency proofs. Additionally, we present the witness and the compiled circuit size for temporal consistency proofs separately in Table V.
In Table I, we can observe that for the case of temporal consistency, the proof size as well as proving and verification key sizes are similar for both datasets. On the other hand, in Table II, the circuit, proving and verification key size for the statistical consistency proofs were constant as expected and minor deviations are observed for the proof and witness sizes across both datasets.
However, we observe an interesting trend with respect to witness and circuit sizes as represented in Table V. It can be seen that the circuit and witness sizes for both datasets exhibit a drop from D values of 1 to 4 before gradually increasing for more extreme D values such as 8, 16 and 32. This behavior can be potentially explained on the basis of row domain padding used by Halo2 proving systems [44] (such as ezkl) which pads rows to the nearest value of $2^{k}$. This behavior is also evidenced by the fact that the witness and circuit sizes almost double from 4 to 12. As a consequence, we can conclude that D=4 might correspond to a sweet spot for the non linear Kalman Filter LSTM based circuit leading to a lower circuit and witness size.
Analyses of Circuit Setup Costs: In Tables III and IV, we present the time incurred during the setup phase of zkSTAR pertaining to model export, settings generation, circuit compilation and circuit setup for HAI and ORNL-PS datasets. In each of these tables we see an almost identical trend as those seen in Table V for each category of setup steps. These trends can be explained on the basis of the fact that the setup of zk-SNARKs relies significantly on the size of the circuits themselves. Therefore, the trends in Tables III and IV can be a natural consequence of the row domain padding adopted by Halo 2 proving backends. This ultimately results in trends wherein a decrease in time consumption is observed from D value of 1 to 4 followed by an increase for more extreme values like 8 and 16.
Computational Analysis of Proving System: We analyze the computational performance for temporal and statistical consistency proof generation for both datasets for varying D values. Our results are presented in Tables VI and VII for HAI and ORNL-PS datasets respectively. The data for these experiments are obtained from multiple runs of varying PSF values and represent the mean and standard deviation for one proof invocation of the circuit. We also provide a speed up value that represents the average case speedup that can be expected when sequentially generating proofs for the entire horizon. We provide the speedup formula and its associated explanation in Equation (58) and Appendix A-A respectively.
The temporal consistency proof generation times are represented by the mean and standard deviation values for D values of 1,4,8 and 16. The proving times are in general stable with minor variations observed with varying D values. We observe that the mean proof generation time decreases slightly from D values of 1 to 8, before increasing slightly at 16. This trend is consistently observed across both datasets for all PSF values. The non-monotonous trends depicted in Tables VI and VII can be due to the non-linear proving costs afforded by ezkl backend. Fundamentally, this trend can be explained on the basis of Table V which highlights the non monotonous behavior of witness and circuit sizes resulting from row domain padding used by Halo2 proving systems. Since the proving times are directly proportional to the circuit size, the trends in Tables VI and VII are exactly aligned with those depicted in Table V.
Lastly, the hypothesis test columns in Tables VI and VII reflect a relatively stable values of average proving time corresponding to statistical consistency proofs. These values are slightly higher for HAI compared to ORNL-PS which might be explained on the basis of the difference in dimensionality between both problems.
Figure 9: Witness Generation Performance with varying D for HAI dataset
Figure 10: Prover Performance for Temporal Consistency with varying D for HAI dataset
System Performance of Witness and Proof Generation: Figures 9 and 10 present trends pertaining to the CPU utilization and memory consumption for only the HAI dataset. we provide the corresponding trends for ORNL-PS in Appendix A-B due to space constraints as well as similar nature of both curves. From Figure 9, we can see greater volatility in CPU utilization and memory consumption for D value of 1. We notice that this trend gradually stabilizes for D values of 8 and 16. The underlying reason for this trend is that proof generation is an inherently lower values of D demand more frequent proving steps that results in higher volatility in CPU utilization and memory consumption. As D grows to 8 and 16, the proving time stabilizes on account of less frequent proving steps encountered during the simulation horizon.
Additionally, we present the proving time system performance using CPU utilization and memory consumption for a random proving window for varying D values. We observe a steady uptick in CPU utilization and memory consumption reaching a value of 1600% and 55GB respectively. The CPU utilization trends reflect the multi-processing capabilities on the hardware that was used which comprised of 16 vCPUs. On the other hand, the memory consumption of 55GB indicates that the proving process consumed more than half the total memory available on the device.
VII Conclusion
In this paper, we address the problem of regulatory compliance in data-driven cyberattack detection for utility stakeholders within large-scale Critical Infrastructure Networks (CINs). Privacy and efficiency constraints make full-fledged ICS data audits impractical, creating a key challenge for regulators seeking trustworthy verification of local detection outcomes. The absence of private, verifiable mechanisms can degrade network-wide situational awareness and increase false positives, resulting in unnecessary downtimes.
To overcome these limitations, we propose zkSTAR, a zkSNARK-based framework that enables publicly verifiable attack detection while preserving complete zero knowledge of local ICS state-space dynamics. zkSTAR provides formal guarantees of temporal and statistical consistency for utility-level detections, linking proof verification directly to the authenticity of attack alarms. For scalability, the framework decomposes proof generation into distinct zkSNARK-based kernels, ensuring computational efficiency and modularity.
We evaluate zkSTAR on a containerized testbed using real-world datasets from HAI and ORNL-PS, demonstrating robust performance across varying temporal consistency intervals and precision scales. System-level analysis further confirms the framework’s stability in terms of CPU and memory utilization. Overall, our results show that zkSTAR enables secure, privacy-preserving, and publicly verifiable cyberattack detection suitable for diverse CIN deployment scenarios.
Appendix A Additional Experimental Results
A-A Proof Setup and Generation Times
The speedup relation used in Tables VI and VII can be given as:
[…]
Dataset DVAL CPU (%) Memory (GB)
HAI 16 1081.584(604.146) 30.135(9.980)
ORNL-PS 16 1081.584(604.146) 30.135(9.980)
Table VIII: CPU and Memory Utilization (mean(std)) for Statistical consistency
Figure 12: Witness Generation Performance with varying D for ORNL-PS dataset
Figure 13: Prover Performance for Temporal Consistency with varying D for ORNL-PS dataset
The speedup computation is based on obtaining the ratio between the average time consumed for the entire set of D unrolled loop iterations when $D=1$ and the corresponding values when $D\neq 1$. Since the time horizon appears in the numerator and denominator, they cancel each other resulting in Equation (58).
A-B Analysis of System Performance of ORNL-PS Dataset
Figures 12 and 13 present the system performance characterized by CPU utilization and memory consumption for ORNL-PS dataset for witness generation and proof generation mechanism.
In Table VIII, we provide the CPU utilization and memory consumption for the statistical consistency proofs for HAI and ORNL-PS datasets.

### v2

@@
 Our approach builds on established residual-based statistical hypothesis testing methods applied to state-space detection models.
-Specifically, we design a two-pronged zk-SNARK architecture that enforces (i) temporal consistency of the state-space dynamics and (ii) statistical consistency of the detection tests—allowing regulators to temporally verify alarm correctness without visibility into utility-level data.
+Specifically, we design a two-pronged zk-SNARK architecture that enforces (i) temporal consistency of the state-space dynamics and (ii) statistical consistency of the detection tests, enabling regulators to verify correctness and prevent suppression of alarms without visibility into utility-level data.
 We formally analyze the soundness and zero-knowledge properties of our framework and validate its practical feasibility through computational experiments on real-world ICS datasets.
 As a result, our work demonstrates a scalable, privacy-preserving alternative for regulatory compliance for ICS driven critical infrastructure networks.
-I Introduction
-Critical Infrastructure Networks (CINs)—such as gas, water, and power systems—are large-scale cyber-physical systems managed by multiple stakeholder entities.
-The growing integration of information and operational technologies (IT-OT) driven by increased IoT adoption has introduced new vulnerabilities to data-driven cyberattacks.
-Such attacks often target the Industrial Control Systems (ICS) of utility stakeholders, aiming to degrade asset performance and potentially trigger cascading failures across the entire network.
-Although attack detection is performed locally at stakeholder ICSs, regulatory entities such as Information Sharing and Analysis Centers (ISACs) aggregate alarms and statistics from all utilities in real time to enhance situational awareness, reduce false positives, and expedite decision-making.
-The effectiveness of ISACs in fulfilling their objectives fundamentally depends on the quality of local ICS detection metrics reported by individual utilities.
-Inconsistent or low-quality local insights can undermine the ISAC’s ability to construct a reliable network-wide situational awareness model, potentially causing unnecessary service disruptions.
-Consequently, there remains a critical need for methods that allow regulators to publicly verify local ICS detection outcomes without resorting to exhaustive utility data audits, which are costly, impractical for real-time operations, and intrusive to utility privacy.
-Therefore, this paper introduces zkSTAR, a framework that enables publicly verifiable detection of data-driven ICS attacks while preserving complete zero knowledge of local ICS state-space dynamics.
-The ICS serves as the primary OT backbone for any CIN utility capable of providing control, actuation and data storage capabilities of infrastructure assets in a scalable manner.
-Data-driven ICS attack detection methods primarily utilize models that characterize the state space of the local utility network which can help unveil deviations from expected steady state conditions using the observed sensor measurements [1].
-These deviations or residuals can then be used as inputs for statistical hypothesis tests to distinguish between normal and abnormal operation profiles subject to Type-I and Type-II error parameters of the test [2].
-An alarm signifying an attack event is triggered locally when the test reject its null hypothesis indicating statistical inconsistencies between the observed and the expected sensor measurements subject to state space dynamics.
-With existing state of the art methods, the only way to conduct regulatory verification of detection outcomes of utility-based state space models is through a full-fledged data audit of ICS operations.
-For such an audit, the utility would have to relinquish the ownership of their local IoT datasets to a regulator raising significant privacy and data governance hurdles.
-Next, utility datasets must be shown to be temporally consistent across the time domain.
-In other words, the regulator must be able to exhaustively certify the state space estimates of successive time steps have been generated from previous estimates.
-Finally, the regulator must also be able to demonstrate statistical consistency which implies that the correct statistical testing procedure was followed to raise alarms.
-As a result, regulatory data audits can become extremely expensive and would come at the cost of significant privacy issues for utilities.
-Therefore, in this paper, we develop zkSTAR, a framework that is geared towards guaranteeing temporal and statistical consistency for state space driven statistical detection of data driven ICS attacks.
-At its core, zkSTAR is specifically designed to demonstrate temporal and statistical consistency of ICS systems with highly non-linear state space dynamics in complete zero knowledge.
-zkSTAR is based on the zkSNARK (zero-knowledge succinct, non-interactive argument of knowledge) paradigm and provides the ability to publicly verify attack detection outcomes for a state space modeling paradigm in a trustworthy fashion.
+1 Introduction
+The convergence of information and operational technologies (IT–OT) through widespread IoT adoption has expanded the attack surface for data-driven cyber intrusions.
+Such attacks often target industrial control systems (ICSs), degrading asset performance and triggering cascading failures across critical infrastructure networks (CINs).
+While detection occurs locally at utilities, regulatory bodies such as Information Sharing and Analysis Centers (ISACs) aggregate alarms in real time to enhance situational awareness and reduce false positives [dhs_oig2].
+The effectiveness of these regulators, however, hinges on the integrity and reliability of local detection data.
+Recent real-world incidents [langner2011stuxnet, jeffries2022cyber, salazar2024tale] have demonstrated that adversaries can compromise both IT and OT layers to obscure, delay, or disable detection of underlying anomalies.
+Such incidents expose a fundamental regulatory blind spot pertaining to verifying the integrity of reported detections without re-auditing raw operational data.
+Moreover, inconsistent or low-quality local insights can undermine the ISAC’s ability to construct a reliable network-wide situational awareness model, potentially causing unnecessary service disruptions [dhs_oig2].
+To address this gap, we introduce zkSTAR, a framework enabling publicly verifiable detection of data-driven ICS attacks in complete zero knowledge of local ICS state-space dynamics.
+zkSTAR addresses a critical regulatory gap by enabling provable temporal and statistical consistency verification of ICS detections.
+Leveraging knowledge-soundness and provably consistent outcomes, it ensures that compromised utilities cannot falsify or suppress alarms even if their IT–OT stacks are breached.
+The framework enables verifiable regulatory oversight while preserving utility data confidentiality [dhs_oig2], eliminating the need for costly, privacy-intrusive audits that have long impeded secure information sharing [nolan2015cybersecurity].
+Additionally, by adopting an on-demand, lazy proof generation paradigm, zkSTAR ensures a low utility overhead as well.
+Conventional ICS anomaly detection relies on state-space models that estimate expected sensor behavior and compute residuals to flag deviations from steady-state conditions [li2020detection].
+Statistical hypothesis tests on these residuals distinguish normal from abnormal operation, triggering alarms when observed measurements significantly diverge from model predictions [li2020deep].
+With existing approaches, regulators can verify such detection outcomes only through full data audits of ICS operations—requiring utilities to relinquish proprietary IoT datasets and raising severe privacy and data-governance concerns.
+Moreover, utilities must demonstrate both temporal consistency wherein successive state estimates are derived from prior ones; and statistical consistency in which alarms follow valid hypothesis test procedures.
+These audits are costly, slow, and incompatible with real-time oversight.
+To overcome these limitations, zkSTAR delivers provable temporal and statistical consistency guarantees for state space driven statistical detection of data driven ICS attacks.
+At its core, zkSTAR is specifically designed to validate consistency claims for ICS systems with highly non-linear state space dynamics in complete zero knowledge.
+zkSTAR is based on the zkSNARK (zero-knowledge succinct, non-interactive argument of knowledge) paradigm providing publicly verifiable attack detection outcomes for a state space modeling paradigm in a trustworthy fashion.
+zkSTAR operates in an on-demand compliance mode, where regulators issue proof-generation requests to utilities whenever verification of local detections is required.
 We present a detailed overview of our contributions as follows:
   • We develop a zkSNARK based prediction framework that leverages an extended Kalman filter based model to detect data driven attacks on ICSs of utility stakeholders in complete zero-knowledge of local sensor or state space dynamics.
-  • We establish comprehensive theoretical guarantees pertaining to the necessary and sufficient conditions for asserting temporally and statistically consistent regulatory compliance in the context of time-series anomaly detection using non-linear state space models.
-  • We design a zkSNARK based kernel decomposition scheme that provides the ability to prove temporal and statistical consistency claims in a computationally efficient manner.
-  • We provide a detailed container-based computational testbed for evaluating the performance of zkSTAR with respect to a detailed case studies involving real-world datasets.
-II Related Works
-Attacks on ICSs that are solely targeted towards the IT systems can be detected and isolated using network traffic monitoring paradigms [3, 4, 5].
-However, data-driven attacks are harder to detect through such mechanisms especially in the ICS context owing to their ability to impact critical pathways in both the IT and OT layers [6] potentially causing critical infrastructure damage.
-Such attacks typically involve manipulation of sensor data so as to effect cascading failures across networked ICS in addition to degrading the performance of critical infrastructure assets.
-Despite robust network traffic monitoring frameworks, such attacks can go undetected since they involve manipulation at the OT layer[7, 8].
-Therefore, several model-based mechanisms have been proposed to detect ICS data-driven attacks [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21].
-Among these, the Kalman filter based mechanism of detecting attack proves to be a robust and flexible option [2, 1, 22, 23, 24] owing to its reliance on residuals emanating from a well-formed state space model.
-Such techniques utilize the residuals obtained from Kalman model predictions and the observed sensor measurements as fundamental inputs to statistical hypothesis tests in order to flag alarms.
-As a result, a Kalman filter based state space estimation mechanism can help detect anomalies emanating from data driven attacks distinguish between routine faults and attacks using degradation models [22] and use decentralized methodologies to raise network-wide alarms [23].
-From the regulatory perspective however, it is essential to not only receive timely notifications regarding attack detection at utility stakeholders but also ensure the integrity of attack detection so as to limit its network-wide impacts and reduce false alarm rates.
-Given the trepidation of utilities in being able to share data with regulatory agencies due to privacy and logistical hurdles, enabling regulatory compliance of local detection frameworks acquires an entirely new degree of complexity.
-As a result, it becomes important to ensure that the right sequence of steps has been followed to determine these insights without belying the underlying data.
-Zero-knowledge proofs form a viable alternative to solve this conundrum with the added capability of complementing privacy mechanisms as well [25].
-Zero-knowledge proofs were originally proposed by Goldwasser et al.
-[26] in which a Prover attempts to convince a Verifier that some mathematical statement is true.
-The basic premise was that the Verifier must not glean any other information about the secrets held by the Prover that validate a set of mathematical statements.
-Original propositions of zero-knowledge proofs were interactive, i.e.
-the Verifier and the Prover exchange a series of messages in order to establish the validity of the mathematical truths.
-However, Blum et al.
-[27] showed that zero-knowledge proofs can also be made non-interactive by using a common short random string (CRS).
-These initial works were followed by establishing the concepts of succinctness that enables efficient verification with minimal overhead [28] which were combined with non-interactivity [29].
-In the past decade, there have been several breakthroughs that have paved the way for practical, computationally feasible implementations of zero-knowledge Succinct, Non-Interactive Arguments of Knowledge (zk-SNARKs) [30, 31, 32, 33, 34].
-Within the domain of machine learning, zero-knowledge proofs have been used to establish the integrity of training [35, 36, 37] and prove model-ownership [38].
-However, there exists a critical gap pertaining to the development of zkSNARKs for time-series anomaly detection using state space dynamics of utility ICS.
-As a result, considering state of the art methods, providing trustworthy data driven attack detection for fulfilling regulatory compliance requirements in complete zero-knowledge in currently infeasible.
-In this paper, we develop zkSTAR, a framework that can generate computationally efficient zkSNARKs of data driven attack detection frameworks that are based on Kalman filter driven utility ICS state space models.
-zkSTAR is primarily geared towards ensuring regulatory compliance in data driven attack detection with full privacy preservation of utility stakeholder data.
-In fact, zkSTAR is capable of enabling regulatory agencies to verify utility level attack detection mechanisms in zero knowledge of their local state space model altogether.
-III Nonlinear State Space Modeling
-For developing zkSTAR, we leverage the extended Kalman Filter state space models owing to their flexibility in handling highly non-linear spatial and temporal interdependencies among several ICS components [39].
+  • We establish formal security guarantees for zkSNARK driven regulatory compliance with temporal and statistical consistency proofs of detection outcomes that also ensure resistance to attack-suppression under knowledge soundness property.
+  • We design a zkSNARK based kernel decomposition scheme that provides the ability to prove temporal and statistical consistency claims in a scalable and computationally efficient manner.
+  • We detail the performance of zkSTAR with respect to detailed case studies involving real-world datasets using a REST-API driven container-based computational testbed.
+Overall, we show that zkSTAR enforces integrity of detection proofs under adversarial falsification, preventing utilities from reusing stale states or forging alarms—making it a security enforcement layer for regulatory trust.
+2 Related Works
+Attacks on industrial control systems (ICSs) that target IT layers can often be detected through network traffic monitoring [caselli2016specification, ye2004robustness].
+However, data-driven attacks that manipulate sensor or control data can evade such approaches by simultaneously impacting both IT and OT layers [urbina2016limiting, kang2014cyber, drias2015taxonomy], potentially causing cascading failures across critical infrastructure.
+More recently, there have been several attack vulnerabilities that have been discovered such as like Industroyer and Industroyer2 [salazar2024tale], Triton [jeffries2022cyber] and Stuxnet [langner2011stuxnet].
+These vulnerabilities have demonstrated the feasibility of joint manipulation of control systems and monitoring data which have motivated the need for robust verifiable detection mechanisms for ICS.
+To address some of these challenges, several model-based mechanisms have been proposed [huang2018online, hoehn2016detection, rahman2017multi].
+Among these, Kalman filter based approaches are particularly robust and flexible [li2020deep, li2020detection, li2021degradation, ramanan2021blockchain, li2022online], leveraging residuals from state-space models as statistical evidence for anomaly detection.
+These residual-based tests differentiate routine faults from malicious attacks [li2021degradation] and can support decentralized, network-wide alarm dissemination [ramanan2021blockchain].
+From the compliance perspective, however, it is essential for the regulator to receive timely notifications regarding attack detection at utility stakeholders in addition to verifying the integrity of attack detection so as to limit its network-wide impacts and reduce false alarm rates [ramanan2021blockchain, dhs_oig2].
+Therefore, regulatory compliance goals would ideally enable stakeholders to provide publicly verifiable detection outcomes that are accurate and trustworthy without exposing private operational data.
+Achieving such verifiable compliance under privacy constraints motivates the use of zero-knowledge proofs [narayan2015verifiable], enabling a prover to convince a verifier of correctness without revealing sensitive information.
+Originating from foundational work by Goldwasser et al.
+[goldwasser2019knowledge] and Blum et al.
+[blum2019non], zero-knowledge proofs evolved into succinct, non-interactive forms (zk-SNARKs) that support efficient verification [kilian1992note, micali2000computationally, gennaro2013quadratic, parno2016pinocchio, ben2013snarks, costello2015geppetto, ben2014succinct].
+Recent applications in machine learning verify training integrity and model ownership [huang2022zkmlaas, wang2022ezdps, nguyen2022preserving, yang2023fedzkp].
+While existing zk-SNARK frameworks enable general-purpose verifiable computation, none target the verification of dynamical system integrity or hypothesis-test correctness which are cornerstones of industrial attack detection.
+Recent works on verifiable differential privacy [narayan2015verifiable] and secure multiparty analytics [froelicher2020drynx] focus primarily on data confidentiality but not on verifiable detection correctness, a gap zkSTAR directly addresses.
+To alleviate this core critical gap, this paper introduces zkSTAR, a framework that enables provable, privacy-preserving verification of Kalman filter–based detection mechanisms, paving the way for regulatory compliance in CINs.
+3 Nonlinear State Space Modeling
+For developing zkSTAR, we leverage the extended Kalman Filter (EKF) state space models owing to their flexibility in handling highly non-linear spatial and temporal interdependencies among several ICS components [coskun2017long].
 Therefore, we characterize the EKF-based, utility-level attack detection model as a foundational component of zkSTAR for achieving regulatory compliance under zero knowledge.
 Within zkSTAR, we develop a state-space formulation capable of identifying attack-induced anomalies through locally executed statistical hypothesis tests.
-III-A Extended Kalman Filter Model
-We present a brief overview of the Extended Kalman Filter model in this section while a detailed description of the model has been provided in Appendix -A.
+3.1 Extended Kalman Filter Model
+We present a brief overview of the EKF model in this section while a detailed description of the model has been provided in Appendix A.1.
 Let g,h{g},{h} denote the state transition and observation functions respectively, $x_{t}\in R^{m}$ represents the latent space embedding, $u_{t}\in R^{m}$ represents the control action and $y_{t}\in R^{d}$ represents noisy sensor measurements from asset sensors.
@@
 $P_{t|t-1}=G_{t}P_{t|t-1}G^{T}_{t}+Q_{t-1},$ (3)
-$S_{t}=H_{t}P_{t|t-1}H^{T}_{t}+R_{t},$ (4)
-$K_{t}=P_{t|t-1}H^{T}_{t}S_{t}$ (5)
+$S_{t}=(H_{t}P_{t|t-1}H^{T}_{t}+R_{t}),$ (4)
+$K_{t}=P_{t|t-1}H^{T}_{t}S_{t}^{-1}$ (5)
 $\hat{x}_{t|t}=\hat{x}_{t|t-1}+K_{t}r_{t}$ (6)
@@
 The state transition and the observation matrices at $t$ given by $G_{t},H_{t}$ respectively are computed using the Jacobians of $g,h$.
-III-B Hypothesis Test for Anomaly Detection
-We consider $r_{t}$ as our residual and $S_{t}=U_{t}\Sigma_{t}U_{t}^{T}$ as the residual covariance.
-We denote the standardized vector of PC scores as $\tau_{t}=(S_{t})^{-1/2}(r_{t})=U_{t}\Sigma_{t}^{-1/2}(r_{t})$ such that $\tau_{t}\sim N(0,I)$ [1, 40].
-This implies that $\tau_{t}||^{2}_{2}\sim\chi^{2}_{p}$, where $p<d$ is the degrees of freedom corresponding to the number of principal components used [40].
+3.2 Hypothesis Test for Anomaly Detection
+We consider residual covariance $S_{t}=U_{t}\Sigma_{t}U_{t}^{T}$ coupled with residual measurements $r_{t}$ for detecting covert data-driven attacks [li2020detection].
+Without loss of generality we note that it is also possible to apply our techniques for detecting false data injection by monitoring $R_{t}$.
+We denote the standardized vector of PC scores as $\tau_{t}=(S_{t})^{-1/2}(r_{t})=U_{t}\Sigma_{t}^{-1/2}(r_{t})$ such that $\tau_{t}\sim N(0,I)$ [li2020detection, hwang2017chi].
+This implies that $\tau_{t}||^{2}_{2}\sim\chi^{2}_{p}$, where $p<d$ is the degrees of freedom corresponding to the number of principal components used [hwang2017chi].
 $\rho=\begin{cases}1,&\text{if }T_{\chi^{2},t}=\|\tau_{t}\|^{2}_{2}>\chi^{2}_{m,\alpha},\\ 0,&\text{otherwise.}\end{cases}$ (8)
-A generic hypothesis testing procedure applicable to sensor and residual monitoring has been discussed in greater detail in Appendix -B.
+A generic hypothesis testing procedure applicable to sensor and residual monitoring has been discussed in greater detail in [li2021degradation].
 We define a measure of probabilistic similarity to help establish the distinctness of two sensor measurement vectors as stated in Definition 1.
@@
 We now proceed towards theoretical characterization of the zkSNARKs to enforce temporal and statistically consistent detection outcomes.
-IV Regulatory Compliance using zkSNARKs
+4 Regulatory Compliance using zkSNARKs
 Detecting attacks requires monitoring covariance matrices of the residuals to drive the underlying $\chi^{2}$ hypothesis tests.
@@
 Therefore, in this paper, we employ zero knowledge proofs to enable regulatory compliance and verification.
-We concentrate on a utility-oriented subsystem represented by a state-space model as outlined in Section III.
-Our privacy considerations specifically address the disclosure risks linked to the public revelation of parameters $C$ and residuals $r_{t}$ derived from the Kalman Filter-based state-space estimation, which are crucial for validating the alarm value $\rho_{t}$.
-As a result, our methodology specifically leverages zk-SNARKs to help assert temporal consistency and statistical consistency for utility-level $\chi^{2}$ hypothesis tests without the need to divulge state space variables.
-IV-A Zero Knowledge Preliminaries
+We concentrate on a utility-oriented subsystem represented by a state-space model as outlined in Section 3 that is meant to be captured in a zkSNARK framework.
+As a result, our methodology specifically leverages zk-SNARKs to assert temporal consistency and statistical consistency for utility-level $\chi^{2}$ hypothesis tests in zero-knowledge of local state space dynamics.
+4.1 Zero Knowledge Preliminaries
 We begin our discussion of zk-SNARK based regulatory compliance schemes for cyber attack detection by establishing four fundamental definitions.
@@
 Definition 2 portrays a hashing function that will be utilized to ensure the integrity of state space estimates.
-We assume that the function hash preserves integrity, one-wayness, hiding and collision resistance properties [41].
+We assume that the function hash preserves integrity, one-wayness, hiding and collision resistance properties [menezes1996handbook].
 Definition 3 (Pretrained State Space Model).
@@
 $\mathcal{M}(\Delta^{in}_{t};\theta)\mapsto\Delta^{out}_{t}$
-$\Delta^{in}_{t}=(\mathcal{H}^{in}_{t},\hat{x}_{t|t},y_{t}),$
+$\Delta^{in}_{t}=(\mathcal{H}^{in}_{t},\hat{x}_{t|t},y_{t})$
 $\Delta^{out}_{t}=(\mathcal{H}^{out}_{t},\rho_{t},\hat{x}_{t+1|t+1},r_{t},K_{t},Q_{t},R_{t})$
@@
 Definition 5 (zk-SNARK Proving Phase).
-The zk-SNARK proof generation phase can be defined by function Prove such that
+The zk-SNARK proving phase is defined by function Prove such that
 […]
@@
 The Prove function consumes the proving key $pk$, $\mathcal{M}$ parametrized by $\theta$ as well as the inputs $\Delta^{in}_{t}$ and outputs $\Delta^{out}_{t}$ at time $t$ to generate a proof artifact $\Pi_{t}$.
-The Prove function needs to be executed for each timestep $t$ at the utility level.
+The Prove function is executed for each timestep $t$ at the utility level.
 Definition 6 (zk-SNARK Verification Phase).
@@
 The Verify function provides output $\Phi_{t}\in\{0,1\}$ which is indicative of successful verification.
-Zero knowledge regulatory compliance for attack detection relies on two fundamental operational attributes of each utility.
-These fundamental operational attributes pertain to temporal consistency and statistical consistency.
+Zero knowledge regulatory compliance for attack detection relies on two fundamental operational attributes of temporal and statistical consistency.
 We argue that these two fundamental operational attributes are necessary and sufficient conditions to enable zero-knowledge regulatory compliance verification.
@@
 In the following subsections, we formally state each operational attribute separately as a precursor for guaranteeing integrity of detection outcomes.
-IV-B Temporal Consistency
-Temporal consistency pertains to the recursive relationship between a posteriori state space estimates at $t$ and $t+1$ respectively.
+4.2 Temporal Consistency
+Temporal consistency (TC) pertains to the recursive relationship between a posteriori state space estimates at $t$ and $t+1$ respectively.
 Temporal consistency constraint encoded in the non-linear Kalman filter model ensures that the posteriori estimate $\hat{x}_{t+1|t+1}$ is a derivative of $\hat{x}_{t|t}$.
@@
 Definition 7 (Temporal Consistency).
-The state space system is considered to be temporally consistent if for each consecutive time step pairs $t,t+1$, the following condition holds.
+The state space system is considered to be temporally consistent if for each consecutive time step pairs $t,t+1$, the following holds.
 $x^{out}_{t}=x^{in}_{t+1}$
@@
 $x^{out}_{t}=\hat{x}_{t|t-1}+K_{t}r_{t},\text{ and, }\hat{x}_{t+1|t}=g(x^{in}_{t+1},u_{t})$
-As a consequence of Definition 7 we can make the temporal consistency claim presented in Theorem IV.1.
-Theorem IV.1 (zk-STAR Temporal Consistency Conditions).
+As a consequence of Definition 7 we can make the temporal consistency claim presented in Theorem 4.1.
+Theorem 4.1 (zk-STAR TC Conditions).
 Given every consecutive time step pairs $t,t+1$ such that $vk_{t}=vk_{t+1}$, $pk_{t}=pk_{t+1}$, temporal consistency can be asserted if and only if the following conditions hold
@@
 Proof.
-Proof provided in Appendix -C
-Theorem IV.1 shows that the integrity of temporal posteriori predictions from an EKF driven state space system can be established using a zkSNARK proof artifact if and only if the verification keys remain invariant; the hash of the inputs to the EKF model at each time step is equal to the hash of the previous EKF output; and the proof artifact itself can be verified using the verification keys.
-In other words, Theorem IV.1 demonstrates that if temporally inconsistent values for carrying out the EKF predictions, then either the verification fails for at least one time step or the keys used for verifying the proof artifact will differ.
-Therefore, to establish temporally consistency from a regulatory perspective, it is necessary and sufficient to check for the invariance of keys, alignment of inputs and outputs and the verification status across successive time steps.
-IV-C Statistical Consistency
-Consider the local attack detection framework characterized by the $\chi^{2}$ hypothesis test presented in Section III-B with a Type-I and Type-II error rates of $\alpha$ and $\beta$ respectively.
+Proof provided in Appendix B.1
+Theorem 4.1 shows that the integrity of temporal posteriori predictions from an EKF driven state space system can be established using a zkSNARK proof artifact if and only if the verification keys remain invariant; the hash commitments of the inputs to the EKF model at each time step is equal to that of the previous EKF output; and the proof artifact itself can be verified using the verification keys.
+In other words, Theorem 4.1 demonstrates that if temporally inconsistent values are used for carrying out the EKF predictions, then either the verification fails for at least one time step or the keys used for verifying the proof artifact will differ.
+Therefore, to establish temporal consistency from a regulatory perspective, it is necessary and sufficient to check for the invariance of keys, alignment of hash commitments and the verification status across successive time steps.
+4.3 Statistical Consistency
+We consider the local attack detection framework characterized by the $\chi^{2}$ hypothesis test presented in Section 3.2.
 The $\chi^{2}$ test relies on the residual and state space estimates provided by $\mathcal{M}$ that is parametrized by $\theta$.
 As a result of Equation (8), the local attack detection framework is expected to yield an alarm value denoted by $\rho_{t}=1$, when an attack is taking place while maintaining $\rho_{t}=0$ under normal circumstances.
-We utilize these fundamental principles to establish the notion of statistical consistency.
+We utilize these fundamental principles to establish the notion of statistical consistency (SC).
 For the purposes of statistical consistency, we denote the genuine and the reported outcomes of the $\chi^{2}$ test as $\rho^{real}_{t}\in\{0,1\}$ and $\rho^{rep}_{t}\in\{0,1\}$ respectively.
@@
 In such cases, expecting to acquire and store a copy of the sensor data measurement at different levels of the ICS that is numerically identical to the one that was originally recorded by the sensor is unrealistic.
-However given the inherent robustness of hypothesis tests towards robustness, utilizing a sensor data measurement that is similar but not necessarily identical to the originally recorded value is sufficient for establishing statistical consistency.
+However, given the inherent robustness of hypothesis tests, utilizing a sensor measurements that is similar but not necessarily identical to the originally recorded value is sufficient for establishing statistical consistency.
 As a result, we develop Lemma 1 as a means to formally establish the notion of probabilistic similarity of sensor measurements and its connection to statistical consistency of the zkSTAR framework.
 Lemma 1.
-Probabilistically similar sensor measurement realizations $\tilde{y}_{t},y_{t}$ are necessary and sufficient conditions for observing statistically coherent outcomes $\tilde{\rho}_{t},\rho_{t}$.
+Probabilistically similar sensor measurement realizations $\tilde{y}_{t},y_{t}$ are necessary and sufficient conditions for observing statistically consistent outcomes $\tilde{\rho}_{t},\rho_{t}$.
 Proof.
-Proof provided in Appendix -D
+Proof provided in Appendix B.2
 Lemma 1 shows that as long as the sensor signal used to carry out EKF predictions is probabilistically similar to the true value of the sensor signal the zkSTAR framework will continue to deliver statistically consistent alarm detection outcomes.
-Moreover, with the help of Lemma 1, we can also state that a data driven ICS attack orchestration might be able to succeed only when the malicious sensor measurements are probabilistically similar to the originally recorded sensor measurements.
-Practically, such attacks end up being inconsequential [22] since they are unable to effect any material damage to the ICS assets in the first place.
-Therefore, Lemma 1 serves as the basis for Theorem IV.2 which derives an integral link between statistical and temporally consistency on one hand and the key invariance and verification status on the other.
-Theorem IV.2 (zk-SNARK Statistical Consistency Conditions).
-Given consecutive time step pairs $t,t+1$ with $(vk_{t},pk_{t})$ and $(vk_{t+1},pk_{t+1})$ and a temporally consistent system, statistical consistency can be guaranteed if and only if $\Phi_{t+1}=1$
+Moreover, with the help of Lemma 1, we can also state that a data driven ICS attack orchestration might in theory go undetected under the zkSTAR framework when the malicious sensor measurements are probabilistically similar to the originally recorded sensor measurements.
+Practically, such attacks end up being inconsequential [li2021degradation] since they are unable to effect any material damage to the ICS assets.
+Therefore, Lemma 1 serves as the basis for Theorem 4.2 which derives an integral link between key invariance, statistical and temporally consistency.
+Theorem 4.2 (zk-SNARK SC Conditions).
+Given consecutive time step pairs $t,t+1$ with $(vk_{t},pk_{t})$ and $(vk_{t+1},pk_{t+1})$, temporal consistency implies statistical consistency.
 Proof.
-Proof provided in Appendix -E
-Theorem IV.1 shows that alarm values obtained from zkSTAR can be deemed statistically consistent if they share the same proving and verification keys, resulting in a successful verification while adhering to Theorem IV.1.
-As a result, a regulatory agency can reject the veracity of reported alarms from utility stakeholder in spite of successful verification if either of the conditions laid down in Theorem IV.2 are not met.
-Additionally, Theorem IV.2 shows that unsuccessful verification of stakeholder proofs implies a statistical inconsistency in alarm detection despite fulfilling temporal consistency claims.
-Lastly, Theorem IV.2 also establishes that temporal inconsistencies necessarily indicate a statistically inconsistent system regardless of the alarm and verification status.
-V Computational Architecture of zkSTAR
-The algorithmic framework of zkSTAR is based on a temporal state space modeling paradigm that is driven by a non-linear LSTM-based Kalman filter model.
-There are three fundamental challenges that arise when attempting to detect data-driven attacks in ICS in zero knowledge.
-First, asserting temporal consistency between successive state space estimations is challenging due to the highly non-linear and dynamic nature of utility based ICS subsystems.
-Second, ensuring the statistical consistency of the detection mechanism without revealing parameters of the statistical test is difficult.
-Lastly, the nature of high-fidelity streaming sensor datasets possess a fine temporal granularity that leads to an additional layer of complexity when characterizing attacks.
-As a result, the zkSTAR framework is fundamentally designed to ensure temporal consistency (TC) and statistical consistency (SC) in the context of high-fidelity streaming sensor data.
-In order to address the challenge of high fidelity sensor information the zkSTAR framework adopts a temporal decomposition scheme that is specifically designed to cater to TC and SC aspects.
-Therefore, in this section, we provide an in-depth overview of the temporal decomposition scheme before proceeding with a technical discussion and algorithmic review of TC and SC components of zkSTAR.
-V-A Temporal Decomposition Scheme
+Proof provided in Appendix B.3
+Theorem 4.2 shows that alarm values obtained from zkSTAR can be deemed statistically consistent if they share the same proving and verification keys and if TC conditions are met according to 4.1.
+As a result, a regulatory agency can deem reported alarms from utility stakeholder as a violation of SC claims despite successful verification if conditions pertaining to key-invariance or temporal consistency are not met.
+We formalize the security model with the help of Theorems 4.1 and 4.2 in the subsequent section.
+4.4 Security Model and Guarantees
+We prove that zkSTAR satisfies knowledge-soundness, guaranteeing that any attempt to forge or reuse stale state estimates for regulatory verification violates the TC-SC consistency relation, thereby ensuring attack-suppression detection.
+For a given timestep $t$, we denote $\Omega_{t}$ to be the witness for corresponding the state space model $\mathcal{M}$ parametrized by $\theta$.
+Using Theorems 4.1 and 4.2, the temporal-statistical consistency relation $\mathcal{R}_{t}$ can be defined as in Equation (9) from the perspective of a non-attack.
+$\mathcal{R}_{t}\Big[(\Delta_{t}^{in},$ $\Delta_{t}^{out}),\Omega_{t}\Big]=\Big[(\mathcal{M}(\Delta_{t}^{in};\theta)=\Delta_{t}^{out})$ (9)
+$\wedge(\mathcal{H}^{in}_{t+1}=\mathcal{H}^{out}_{t})\wedge(T_{\chi^{2},t}(\tau_{t})\leq\chi^{2}_{m,\alpha})\Big].$
+We consider a utility prover $\mathcal{U}$ undergoing a data-driven cyberattack starting at timestep $t^{attack}>>t^{init}$ where $t^{init}$ represents the initial time step pertaining to the setup phase (Definition 3).
+For concreteness, we base our analysis on the Halo2 backend for establishing knowledge-soundness.
+For the analysis of the security model and to establish knowledge-soundness, we assume that the objective of the malicious attacker is to carry out the attack while suppressing attack detection.
+Moreover, we assume that the compromised utility possesses an extractor $\mathcal{E}^{\mathcal{U}}$ that produces a valid proof $\Pi_{t}$ which results in $\Phi_{t}=1$ with an incorrect witness $\mathcal{E}^{\mathcal{U}}(\Delta_{t}^{in},\Delta_{t}^{out})\neq\Omega_{t}$, with a non-negligible probability.
+Formally, we define attack suppression using Definition 9.
+Definition 9 (Detection Suppression).
+The detection framework characterized by $\mathcal{M}(.;\theta)$ is deemed to have been suppressed when the following conditions hold
+[…]
+$\big((\Delta_{t}^{in},\Delta_{t}^{out}),\,\mathcal{E}^{\mathcal{U}}(\Delta_{t}^{in},\Delta_{t}^{out})\big)\notin\mathcal{R}_{t}$
+As a result using the TC and SC claims detailed in Theorems 4.1 and 4.2, we characterize the infeasibility of detection suppression in Theorem 4.3.
+Theorem 4.3 (Detection Suppression and Knowledge Soundness).
+Given a utility prover $\mathcal{U}$ a timestep $t>t^{attack}$, detection suppression occurs with only negligible probability under the knowledge-soundness property of the underlying proving system.
+Furthermore, infeasibility of detection suppression is equivalent to violation of either proof verification or of the temporal-statistical consistency relation $\mathcal{R}_{t}$.
+Proof.
+Proof provided in Appendix B.4. ∎
+Theorem 4.3 fundamentally demonstrates that a utility undergoing an attack must either fail proof verification or violate TC-SC relations enforced by zkSTAR.
+In other words, an under-attack utility cannot fabricate or reuse stale system states without detection, since any such attempt would contradict the knowledge-soundness of the Halo2 backend rendering attack suppression cryptographically infeasible.
+5 Computational Architecture of zkSTAR
+The zkSTAR architecture is based on the existence of two components comprising the utility as the prover and the regulator as the verifier.
+As a result, the algorithmic core of zkSTAR focuses on the utility component, encapsulating the utility level EKF model discussed in Section 3 in a zero-knowledge framework.
+The regulatory component deals with verifying the utility generated proof artifacts and asserting TC, SC and detection suppression claims of utilities.
+The framework is designed to generate proofs in an on-demand, lazy fashion using keys generated at the utilities.
+As a result, utilities commit to the EKF model a priori and use the keys to generate only witnesses in real time containing state space estimates including residuals, covariances and test statistics for each detection window.
+A regulator can invoke proof generation on-demand at the utility level for specific detection windows as per verification needs.
+When attempting to encapsulate the utility level EKF model into a zkSNARK structure, we face three crucial challenges.
+First, asserting temporal consistency is challenging due to the highly non-linear and dynamic nature of utility based ICS subsystems.
+Second, ensuring statistically consistent detection outcomes without revealing parameters of the statistical test is difficult.
+Lastly, high-fidelity streaming sensor datasets possess a fine temporal granularity that leads to an additional layer of computational complexity when characterizing attacks.
+In order to address these challenges, the zkSTAR framework adopts a temporal decomposition scheme that specifically caters to TC and SC aspects.
+5.1 Temporal Decomposition Scheme
 We consider a real-time attack detection horizon that is discretized into a sequence of detection windows denoted by set $\mathcal{W}_{i},\forall i\in\{1,2,\ldots\}$.
-Each detection window $\mathcal{W}_{i}$ is a set of $W$ time intervals corresponding to utility ICS state space model.
-Each time interval is represented by $w_{ij},\forall j\in\{1,W\}$ and is primarily geared towards enforcing temporal consistency across multiple prediction time steps.
-Every TC interval is denoted by $w_{ij},i\in\{1,2,\ldots\},j\in\{1,W\}$ and consists of $D$ consecutive timestamps such that $w_{ij}=\{t_{w_{ij}},t_{w_{ij}}+1,t_{w_{ij}}+2\ldots t_{w_{ij}}+D\}$, where $t_{w_{ij}}$ is the starting timestamp of detection window $w_{ij}$.
-The relationship between each detection window $\mathcal{W}_{i}$ and its constituent TC intervals $w_{ij}$ can be compactly denoted using the following equations.
+Each detection window $\mathcal{W}_{i}$ is in turn divided into a set of $W$ TC intervals.
+Each TC interval is represented by $w_{ij},\forall j\in\{1,W\}$ and is primarily geared towards enforcing TC across multiple prediction time steps.
+Every TC interval consists of $D$ consecutive timestamps such that $w_{ij}=\{t_{w_{ij}},t_{w_{ij}}+1,t_{w_{ij}}+2\ldots t_{w_{ij}}+D\}$, where $t_{w_{ij}}$ is the starting timestamp of TC interval $w_{ij}$.
+The relationship between each detection window $\mathcal{W}_{i}$, its constituent TC intervals $w_{ij}$ and corresponding timesteps can be compactly denoted using Equations (10) and (11).
 $\mathcal{W}_{i}=\{w_{i1},w_{i2},\ldots w_{iW}\},\ \forall i\in 1,2,\ldots$ (10)
@@
 Without loss of generality, let us consider a detection window $\mathcal{W}$ and its constituent TC intervals $w_{j}\in\mathcal{W},\ j\in\{1,W\}$.
-We know from Section III that for any time step $t_{w}+k\in w$, $r_{t_{w}+k}\sim N(0,S_{t_{w}+k})$ where $r_{t_{w}+k},S_{t_{w}+k}$ are the residual and the covariance matrix predicted by the pre-trained EKF model for time step $t_{w}+k$.
-Since $r_{t_{w}+k}\sim N(0,S_{t_{w}+k})$, the residuals and covariance matrices can be aggregated across all TC intervals for detection window to yield $r_{W}$ and $S_{W}$ as represented in Equation (11).
+We know from Section 3 that for any time step $t_{w}+k\in w$, $r_{t_{w}+k}\sim N(0,S_{t_{w}+k})$ where $r_{t_{w}+k},S_{t_{w}+k}$ are the residual and the covariance matrix predicted by the pre-trained EKF model for time step $t_{w}+k$.
+Since $r_{t_{w}+k}\sim N(0,S_{t_{w}+k})$, the residuals and covariance matrices can be aggregated across all TC intervals in a detection window to yield $r_{W}$ and $S_{W}$ denoted in Equation (12).
 $r_{W}=\sum\limits_{j=1}^{W}\sum\limits_{k=1}^{D}r_{t_{w_{j}}+k}\ \text{,}\ S_{W}=\sum\limits_{j=1}^{W}\sum\limits_{k=1}^{D}S_{t_{w_{j}}+k}$ (12)
 As a result of hierarchical decomposition, there exists a distinct detection alarm $\rho_{W}$ for each detection window $\mathcal{W}$.
-The temporal decomposition scheme paves the way to leverage zkSNARKs that can help enforce integrity of state space predictions based on a pre-committed EKF model.
+The temporal decomposition scheme paves the way for zkSNARKs to help enforce integrity of state space predictions based on a pre-committed EKF model.
 However, the zkSNARK circuit size grows exponentially with the growth in the detection window size making the proof generation step computationally expensive especially for larger systems.
-Therefore, in order to achieve computational efficiency, while retaining the practicality of zkSNARKs, we adopt a kernel based strategy that decouples compute tasks pertaining to an EKF based state space attack detection framework.
-Specifically, we design zkSNARK based kernels that target individual sub tasks pertaining to temporal consistency of EKF state space predictions as well as statistically consistent outcomes of the detection model that operates on batched EKF residual outputs.
-We also note that state of the art zkSNARK frameworks currently do not support linear algebra operations, such as linear solvers, SVD decompositions while lacking the capability of evaluating derivatives of complex sub-modules.
-This lack of support largely emanates from the significant computational cost of conducting these operations in full zero knowledge making them impractical.
-In order to circumvent these constraints, we adopt a compute-offline-validate-online approach wherein computationally expensive steps such as derivative evaluations, linear system solutions and SVD decompositions are carried out off-circuit but are evaluated for consistency with the state space model within the zkSNARK circuit.
+Therefore, in order to balance practicality and computational efficiency, we adopt a kernel based strategy that decomposes various compute tasks pertaining to an EKF model.
+Specifically, we design zkSNARK based kernels that target individual sub tasks pertaining to TC and SC attributes of EKF predictions and detection outcomes capable of operating on batched EKF residual outputs.
+We note that current zkSNARK frameworks such as ezkl lack support for linear solvers, SVD decompositions, and derivative evaluations of complex sub modules due to their high computational cost in zero knowledge [ezkl2024].
+To address these limitations, we employ a compute-offline–validate-online strategy, performing such expensive operations off-circuit while verifying their consistency with the state-space model within the zkSNARK circuit.
 Figure 1: Temporal consistency handling in zkSTAR
-V-B zkSNARK based Temporal Kernels
-For ensuring temporally consistent outcomes, our kernel design strategy involves the breaking down distinct components of the EKF model inference.
+5.2 zkSNARK based Temporal Kernels
+For ensuring temporally consistent outcomes, our kernel design strategy involves breaking down distinct components of the EKF model inference.
 Specifically, we decompose the EKF inference over a single TC interval into distinct phases that pertain to validating the consistency of the Kalman gain matrix with the state covariance and process and measurement Jacobians; inference of the pre-committed EKF model on one timestamp of the TC interval; and accumulating the residual and its covariance sum across the entire TC interval.
-Kalman Reconstruction Check (KRC) Kernel: The EKF model presented in Section III relies on successive linear solves of Equation (28) in order to determine the Kalman gain $K_{t}$.
+Kalman Reconstruction Check (KRC) Kernel: The EKF model presented in Section 3 relies on successive linear solves of Equation (5) in order to determine the Kalman gain $K_{t}$.
 However, this poses a major challenge since implementing linear system solvers within a zkSNARK is computationally very expensive.
 Therefore, we develop a computational kernel tailored for the zkSNARK paradigm that demonstrates that offline computed matrices Kalman gain $K$, LSTM Jacobian $H$, residual covariance $S$ and state covariance $P$ are consistent with respect to the state space EKF model.
-The result is the KRC kernel depicted in Algorithm 1 which yields a boolean output $\kappa_{t}$ depending on whether the error in Equation (28) is below or above threshold $\epsilon_{KC}$
+The result is the KRC kernel depicted in Algorithm 1 which yields a boolean output $\kappa_{t}$ depending on whether the error in Equation (5) is below or above $\epsilon_{KC}$
 Algorithm 1 zkSTAR Kernel for Kalman Gain Reconstruction
-1:function KRC-Kernel($K$, $S^{-1}$, $H$, $P$, $\epsilon$)
+1:function KRC-Kernel($K$, $S$, $H$, $P$, $\epsilon$)
 […]
@@
 4: compute priori covariance $P_{t|t-1}$ using Equation (3)
-5: compute inverse innovation $S_{t}^{-1}$ using Equation (27).
+5: compute innovation $S_{t}$ using Equation (4).
 […]
@@
 Algorithm 2 represents the computational kernel for EKF within the zkSTAR framework denoted by EKF-Kernel.
-For a specific timestamp $t$, EKF-Kernel computes the priori state and covariance estimates $\hat{x}_{t|t-1},P_{t|t-1}$ as well as the resulting residuals $r_{t}$ using Equations (24), (26) and (25) respectively.
-Using Equation (27), the kernel estimates the inverse of the innovation matrix or the residual covariance $S_{t}^{-1}$.
-Next, the KRC-Kernel is invoked to validate the consistency of the Kalman gain matrix $K_{t}$ with respect to the Jacobians $G_{t}$, $H_{t}$ as well as the posteriori covariance $P_{t|t}$.
+For a timestamp $t$, EKF-Kernel computes the priori state and covariance estimates $\hat{x}_{t|t-1},P_{t|t-1}$ as well as the resulting residuals $r_{t}$ using Equations (1), (3) and (2) respectively.
+Using Equation (4), the kernel estimates the innovation or the residual covariance matrix $S_{t}$.
+Next, the KRC-Kernel is invoked to validate the consistency of the Kalman gain matrix $K_{t}$ with respect to the Jacobian $H_{t}$, the residual covariance $S_{t}$, as well as the posteriori covariance $P_{t|t}$.
 Subsequently, the posteriori state and covariance estimates are updated according to Equations (6) and (7).
-The kernel returns the current residual $r_{t}$, the residual covariance $S_{t}$ and the Kalman gain reconstruction status $\kappa_{t}$.
+The kernel returns the current residual $r_{t}$, the residual covariance $S_{t}$ and the Kalman reconstruction status $\kappa_{t}$.
 Temporal Consistency (TC) Kernel: While the EKF kernel can enforce state space dynamics for each timestamp $t$, it cannot independently ensure the integrity of state space estimates across multiple time stamps.
 In order to do so, we need a mechanism to assert the integrity between distinct timestamps.
-Therefore, we develop a temporal consistency (TC) kernel that is geared towards ensuring the integrity across dictinct timestamps.
-Using the temporal decomposition scheme highlighted in Section V-A we focus on a single TC interval denoted by $w$ to aid our discussion.
+Therefore, we develop a temporal consistency (TC) kernel that is geared towards ensuring the integrity across distinct timestamps.
+Using the temporal decomposition scheme highlighted in Section 5.1 we focus on a single TC interval denoted by $w$ to aid our discussion.
 We denote the inputs to the TC kernel in terms of three sets: $\mathcal{X}^{TC}_{w-1},\mathcal{R}^{TC}_{w-1}$ and $\mathcal{K}^{TC}_{w-1}$ which represent state, residual and KRC status estimates from the previous TC interval w1w-1.
@@
 $\mathcal{K}^{TC}_{w-1}=\underbrace{(\kappa_{w-1}||N^{\kappa}_{w-1})}_{\text{nonce appended KRC status}}$ (16)
-Each set in turn consists of the raw values that are appended with a unique randomly chosen none value to provide a good degree of entropy to the resulting hash strings.
+Each set in turn consists of the raw values that are appended with a unique randomly chosen nonce value to provide a robust degree of entropy to the resulting hash strings.
 As a result, $\mathcal{X}^{TC}_{w-1}$ consists of $(x_{w-1|w-1}||N^{x}_{w-1}),(P_{w-1|w-1}||N^{P}_{w-1})$ where $x_{w-1|w-1}$, $P_{w-1|w-1}$, $N^{x}_{w-1}$ and $N^{P}_{w-1}$ denote the posteriori state and covariance estimate and their corresponding nonce values from the previous TC interval.
-Similarly, $\mathcal{R}^{TC}{w-1}$ comprises the residual and covariance pairs $(r{w-1}||N^{r}_{w-1})$ and $(S_{w-1}||N^{S}_{w-1})$, where $r_{w-1}$ and $S_{w-1}$ represent the innovation and its associated covariance from the previous window, and $N^{r}_{w-1}$ and $N^{S}_{w-1}$ are the corresponding nonce values ensuring stochastic uniqueness of their hash representations.
-Finally, $\mathcal{K}^{TC}{w-1}$ encapsulates $(\kappa_{w-1}||N^{\kappa}_{w-1})$, where $\kappa{w-1}$ denotes the binary KRC indicator and $N^{\kappa}_{w-1}$ is its appended nonce value used to prevent deterministic linkage across temporal proofs.
+Similarly, $\mathcal{R}^{TC}_{w-1}$ comprises the residual and covariance pairs $(r{w-1}||N^{r}_{w-1})$ and $(S_{w-1}||N^{S}_{w-1})$, where $r_{w-1}$ and $S_{w-1}$ represent the innovation and its associated covariance from the previous window, and $N^{r}_{w-1}$ and $N^{S}_{w-1}$ are the corresponding nonce values ensuring stochastic uniqueness of their hash representations.
+Finally, $\mathcal{K}^{TC}_{w-1}$ encapsulates $(\kappa_{w-1}||N^{\kappa}_{w-1})$, where $\kappa_{w-1}$ denotes the binary KRC indicator and $N^{\kappa}_{w-1}$ is its appended nonce value used to prevent deterministic linkage across temporal proofs.
 $\mathcal{Z}^{TC}_{t}=\{\underbrace{(y_{t}||N^{y}_{t}),(G_{t}||N^{G}_{t}),(H_{t}||N^{H}_{t}),(K_{t}||N^{K}_{t})}_{\text{nonce appended sensor data, Jacobian and Kalman Gain}}\}$ (17)
@@
 The set $\mathcal{Z}^{TC}_{t}$ contains the nonce-appended sensor measurements and model parameters $(y_{t}||N^{y}_{t})$, $(G_{t}||N^{G}_{t})$, $(H_{t}||N^{H}_{t})$, and $(K_{t}||N^{K}_{t})$ corresponding to the sensor data, the Jacobians, and Kalman gain, respectively.
-Each element in $\mathcal{Z}^{TC}{t}$ is locally randomized through its associated nonce to ensure that the resulting hashed representations remain unlinkable across successive EKF invocations.
+Each element in $\mathcal{Z}^{TC}_{t}$ is locally randomized through its associated nonce to ensure that the resulting hashed representations remain unlinkable across successive EKF invocations.
 Collectively, the windowed set $\mathcal{Z}^{TC}_{w}$ aggregates these per-timestep inputs over the duration of the TC interval $w$, thereby serving as the complete batch of EKF inputs required for evaluating the zkSTAR TC kernel and generating the proof of temporal consistency for that interval.
@@
 […]
-8: accumulate residual sum $r_{w}\leftarrow r_{w}+r_{t}$
-9: accumulate residual covariance sum $S_{w}\leftarrow S_{w}+S_{t}$
-10: accumulate KRC status $\kappa_{w}\leftarrow\kappa_{w}\circ\kappa_{t}$
-11:end for
-12:#Append respective nonce strings from $\mathcal{N}^{TC}_{w}$
+[…]
+[…]
+10: accumulate residual sum $r_{w}\leftarrow r_{w}+r_{t}$
+11: accumulate residual covariance sum $S_{w}\leftarrow S_{w}+S_{t}$
+12: accumulate KRC status $\kappa_{w}\leftarrow\kappa_{w}\circ\kappa_{t}$
+13:end for
+14:#Append nonce strings from $\mathcal{N}^{TC}_{w}$
 […]
@@
 […]
-16:return $\mathcal{X}^{TC}_{w},\mathcal{R}^{TC}_{w},\mathcal{K}^{TC}_{w}$
-17:end function
-Algorithmically, the kernel is represented using the TC-Kernel function presented in in Algorithm 3.
+18:return $\mathcal{X}^{TC}_{w},\mathcal{R}^{TC}_{w},\mathcal{K}^{TC}_{w}$
+19:end function
+Algorithmically, the kernel is represented using the TC-Kernel function presented in Algorithm 3.
 The kernel consumes the outputs of the previous TC interval w1w-1 and iterates through the timestamps $\{t_{w}\ldots t_{w}+D\}$ belonging to the current TC interval $w$.
@@
 The residual, covariance and KRC status outputs of the EKF kernel are accumulated across all the timestamps of the current TC interval $w$ through summation and multiplication operations respectively.
-At each iteration of the kernel, Finally, the kernel returns as the overall outputs of the TC kernel itself.
+At each iteration, the kernel returns the overall outputs pertaining to the state, residual and reconstruction status check for the corresponding TC interval.
 Figure 2: Statistical consistency estimation in zkSTAR
-V-C zkSNARK based Statistical Kernels
-Applying the $\chi^{2}$ test on the accumulated residuals and covariance requires us to carry out SVD decomposition to evaluate the test statistic and the alarm criteria.
-Conducting the SVD decomposition, evaluating the corresponding PC scores, evaluating the test statistic and the subsequent alarm value must be conducted completely in zero knowledge.
+5.3 zkSNARK based Statistical Kernels
+Applying the $\chi^{2}$ test on the accumulated residuals and covariance requires us to carry out SVD decomposition to evaluate the test statistic and the alarm criteria in complete zero-knowledge.
 Therefore, we design a sequence of statistical zkSNARK kernels to help achieve efficiency in zkSTAR.
@@
 Additionally, we compute the SVD of the accumulated covariance matrix in an offline manner and validate its consistency using the SVD reconstruction kernel denoted by SVD-Kernel.
-More specifically, we represent the inputs to the SVD kernel in terms of $\mathcal{Z}^{SVD}_{W}$ as stated in Equation (19) which consists of the singular vectors and the square roots of the singular values.
+We represent the inputs to the SVD kernel in terms of $\mathcal{Z}^{SVD}_{W}$ as stated in Equation (20) which consists of the singular vectors and the square roots of the singular values.
 $\mathcal{Z}^{SVD}_{W}=\{\underbrace{(U_{W}||N^{U}_{W}),(\Sigma^{-\frac{1}{2}}_{W}||N^{\Sigma}_{W})}_{\text{nonce appended SVD values}}\}$ (20)
@@
 […]
-5:#Set UCL and default alarm values for W
+5:#Set UCL and default alarm for W
 […]
@@
 The kernel outputs a tuple of values consisting of the alarm $\rho_{W}$, the SVD status check $\eta_{W}$ as well as the cumulative KCR status check of $\kappa_{W}$.
-V-D zkSTAR Algorithmic Framework
-We describe the overall zkSTAR algorithmic framework in Algorithm 6 that encapsulates the temporal and statistical consistency kernels.
+Algorithm 6 zkSTAR Algorithmic Framework
+1:#Set visibility for TC circuit
+2:#input;output;param visibility
+[…]
+4:
+5:#Setup proving circuit for TC
+[…]
+7:
+8:#Set visibility for SC circuit
+9:#input;output;param visibility
+[…]
+11:
+12:#Setup proving circuit for SC
+[…]
+14:
+15:distribute $vk^{TC},vk^{SC}$ publicly
+16:
+17:for $i=1,2\ldots$ do $\triangleright$ rolling detection windows $\mathcal{W}_{i}$
+18:for $j=1,2,\ldots W$ do $\triangleright$ TC interval $w_{ij}$
+19: generate nonce values $\mathcal{N}^{TC}_{j}$
+20: compute $\mathcal{Z}^{TC}_{j}$ using Equation (18)
+21: compute $\mathcal{X},\mathcal{R},\mathcal{K}]_{j-1}$ using Equations (14)-(16).
+[…]
+[…]
+24:
+25:#Generate TC proof for interval $j$
+[…]
+27:
+28: extract $\mathcal{X},\mathcal{R},\mathcal{K}]_{j}$ from $\mathcal{O}^{TC}_{j}$
+29:end for
+30: extract $S_{W}$ from $\mathcal{R}_{W}$
+[…]
+32: generate nonce values $\mathcal{N}^{SVD}$
+33: compute $\mathcal{Z}^{SVD}_{W}$ using SVD values $U_{W},\Sigma_{W}^{-\frac{1}{2}},\mathcal{N}^{SVD}$
+[…]
+35:
+36:#Generate SC proof for window $\mathcal{W}_{i}$
+[…]
+38:
+39:return $\mathcal{P_{W_{i}}}=\Big[\Big(\Pi^{TC}_{w_{i1}}\ldots\Pi^{TC}_{w_{iW}}\Big),\Pi^{SC}_{\mathcal{W_{i}}}\Big]$
+40:end for
+5.4 zkSTAR Algorithmic Framework
+We describe the overall zkSTAR algorithmic framework in Algorithm 6 that encapsulates the TC and SC kernels.
 The framework primarily executes on a rolling horizon of detection windows characterized by $\mathcal{W}_{i}$ with each window consisting of $W$ TC intervals of $D$ timestamps each.
@@
 […]
-Algorithm 6 zkSTAR Algorithmic Framework
-1:#Set visibility for TC circuit
-[…]
-3:
-4:#Setup proving circuit for TC
-[…]
-6:
-7:#Set visibility for SC circuit
-[…]
-9:
-10:#Setup proving circuit for SC
-[…]
-12:
-13:distribute $vk^{TC},vk^{SC}$ publicly
-14:
-15:for $i=1,2\ldots$ do $\triangleright$ detection window $\mathcal{W}_{i}$
-16:for $j=1,2,\ldots W$ do $\triangleright$ TC interval $w_{ij}$
-17: generate nonce values $\mathcal{N}^{TC}_{j}$
-18: compute $\mathcal{Z}^{TC}_{j}$ using Equation (17)
-19: compute $\mathcal{X},\mathcal{R},\mathcal{K}]_{j-1}$ using Equations (13)-(15).
-[…]
-[…]
-22:
-23:#Generate TC proof for interval $j$
-[…]
-25:
-26: extract $\mathcal{X},\mathcal{R},\mathcal{K}]_{j}$ from $\mathcal{O}^{TC}_{j}$
-27:end for
-28: extract $S_{W}$ from $\mathcal{R}_{W}$
-[…]
-30: generate nonce values $\mathcal{N}^{SVD}$
-31: compute $\mathcal{Z}^{SVD}_{W}$ using SVD values $U_{W},\Sigma_{W}^{-\frac{1}{2}},\mathcal{N}^{SVD}$
-[…]
-33:
-34:#Generate SC proof for window $\mathcal{W}_{i}$
-[…]
-36:
-37:return $\mathcal{P_{W_{i}}}=\Big[\Big(\Pi^{TC}_{w_{i1}}\ldots\Pi^{TC}_{w_{iW}}\Big),\Pi^{SC}_{\mathcal{W_{i}}}\Big]$
-38:end for
-For each detection window $\mathcal{W}i$, the algorithm iterates through the TC intervals wijw{ij} of length $W$.
-Within each TC interval, a fresh set of nonce values $\mathcal{N}^{TC}{j}$ is generated to enable efficient hiding for hash outputs.
-Using these nonces, the algorithm constructs the nonce-appended input set $\mathcal{Z}^{TC}{j}$ according to (17), which encapsulates sensor data, model Jacobians, and Kalman gain parameters.
-The posterior state, residual, and kernel reconstruction variables from the previous interval, $\mathcal{X},\mathcal{R},\mathcal{K}]{j-1}$, are then obtained from (13)–(15).
-Together, these form the TC-kernel input $\mathcal{I}^{TC}{j}={[\mathcal{X},\mathcal{R},\mathcal{K}]{j-1},[\mathcal{Z}^{TC},\mathcal{N}^{TC}]{j}}$, which is passed to the TC-Kernel for zkSNARK-based verification of temporal consistency.
+For each detection window $\mathcal{W}i$, the algorithm iterates through the TC intervals $w_{ij}$ of length $W$.
+Within each TC interval, a fresh set of nonce values $\mathcal{N}^{TC}_{j}$ is generated to enable efficient hiding for hash outputs.
+Using these nonces, the algorithm constructs the nonce-appended input set $\mathcal{Z}^{TC}_{j}$ according to Equation (18), which encapsulates sensor data, model Jacobians, and Kalman gain parameters.
+The posterior state, residual, and kernel reconstruction variables from the previous interval, $\mathcal{X},\mathcal{R},\mathcal{K}]_{j-1}$, are then obtained from (14)–(16).
+Together, these form the TC-kernel input $\mathcal{I}^{TC}_{j}={[\mathcal{X},\mathcal{R},\mathcal{K}]_{j-1},[\mathcal{Z}^{TC},\mathcal{N}^{TC}]_{j}}$, which is passed to the TC-Kernel for zkSNARK-based verification of temporal consistency.
 The kernel outputs the updated tuple $\mathcal{O}^{TC}{j}={[\mathcal{X},\mathcal{R},\mathcal{K}]{j}}$, whose elements are extracted and propagated to the next interval.
@@
 This output constitutes the verifiable zkSNARK-backed anomaly indicator used by zkSTAR for secure, decentralized model validation.
-VI Experimental Results
+5.5 Assumptions and Limitations
+The zkSTAR framework relies on several assumptions pertaining to the properties of the zero-knowledge proving system.
+First, since zkSTAR uses the Halo2 proof generation paradigm as a consequence of using the ezkl framework [ezkl2024], it implicitly depends on elliptic-curve cryptographic primitives and on the integrity of the Common Reference String (CRS).
+As a result, we assume that the CRS is generated through a publicly verifiable ceremony per-utility.
+Second, we assume that the EKF model is well-formed and is strongly sensitive to the Jacobian matrices [huang2008analysis] implying that slight deviations in SVD or Kalman gain computations would lead to inconsistent state and residual estimates violating either TC or SC conditions.
+Third, we assume that verifier is honest-but-curious and that public values of the proof artifacts do not violate privacy of utility state space which is a common requirement of existing regulators [doe].
+Finally, we assume that communication channels are assumed to preserve integrity but not confidentiality which can be alleviated by using transport-level encryption like TLS for privacy during proof transmission.
+Under these assumptions, zkSTAR achieves end-to-end regulatory verification with cryptographic soundness and provable privacy, while maintaining computational feasibility for real-world industrial deployments.
+6 Experimental Results
 Figure 3: HAI Dataset: Detection Quality with varying D and fixed PSF
 Figure 4: ORNL-PS Dataset: Detection Quality with varying D and fixed PSF
-Figure 5: HAI Dataset: Detection Quality with varying PSF and fixed D
-Figure 6: ORNL-PS Dataset: Detection Quality with varying PSF and fixed D
-For experimental studies, we utilize the HAI [42] and ORNL-PS [20, 43] dataset.
-To build a asset level anomaly detector, we generate pretrained non-linear models using the EKF approach to accurately reflect the state-space of the critical asset subsystems as described in Section III.
+For experimental studies, we utilize the HAI [hai] and ORNL-PS [pan2015classification, pan2015developing] datasets.
+To build an asset level anomaly detector, we generate pretrained non-linear models using the EKF approach to accurately reflect the state-space of the critical asset subsystems as described in Section 3.
+The training methodology for EKF models is briefly discussed in Appendix A.2.
 We encode the temporal and statistical components of these models in corresponding zk-SNARKs as described in previous sections.
-As a result, the zkSTAR framework yields that encode temporal and statistical consistency with respect to detection outcomes.
+As a result, the zkSTAR framework encodes temporal and statistical consistency with respect to detection outcomes.
 We carry out detection based on a pre-determined value of upper control limit (UCL) that is representative of the levels of significance used for the $\chi 2$ hypothesis test.
@@
 We use a value of 5e35e-3 and 2e22e-2 for HAI and ORNL-PS datasets that were obtained after rigorous empirical studies.
+All of our results include comparison with the non-ZKP version as our benchmark to judge the detection quality.
 System Implementation Details: All experiments were carried out on a virtual machine (VM) running Ubuntu 24.04 with 100GB of RAM and 16 vCPUs using Python 3.11 with the detection model inference using PyTorch 2.7.1.
-We utilize the ezkl library for generating all the zkSNARKs associated with the zkSTAR frameworkWe provide a native as well as a container based mechanism for executing the zkSTAR framework.
-The native method is primarily meant for comprehensive analysis on a high-performance computing (HPC) clusters to evaluate detection accuracy and computational correctness with varying value of residual aggregation window (referred to as the D value) and precision scale factor (PSF value).
-The container based method is primarily meant to evaluate real-world system performance in terms of CPU utilization and memory consumption for the distinct processes of witness generation as well as temporal consistency and statistical consistency proofs.
-We utilize separate containers to represent utility as the prover and regulator as the verifier with an interactive REST API driven interface for evaluation and deployment.
-VI-A Detection Quality of zkSNARK outcomes
-For assessing the detection quality, we carry out two sets of experiments that depict the effects of varying the PSF with a fixed D and vice versa as well.
-Specifically, these experiments pertain to a total detection run-time window of 10 minutes in which the attack starts roughly in the middle.
-Additionally, the attack duration window includes the actual attack occurrence as well as a buffer time of 3 minutes for the HAI dataset.
-For the ORNL-PS dataset, the attack duration window lasts until the end of the detection run-time window of 10 minutes.
-Even though the HAI attack scenario lasts 164 secs, we include a buffer time for HAI due to the highly non-linear nature of the detection model that can induce slight delays before anomalies can acquire statistical significance.
-For these experiments, we consider PSF values of 8,10 and 12, while using 4, 8, 16 and 32 as the D values.
+We utilize the ezkl library [ezkl2024] for generating all the zkSNARKs associated with the zkSTAR framework.
+We implement zkSTAR using both native and containerized execution modes.
+The native mode targets high-performance computing (HPC) clusters for evaluating detection accuracy and computational correctness across varying residual aggregation windows ($D$) and precision scale factors (PSF).
+The containerized mode assesses real-world performance, including CPU and memory usage during witness generation and TC and SC proof generation and verification.
+Separate containers emulate the utility (prover) and regulator (verifier), communicating via REST APIs and managed through the zkstarctl command-line interface.
+6.1 Detection Quality of zkSNARK outcomes
+To evaluate detection quality, we conduct two experiments examining the effects of varying the PSF with fixed $D$ and vice versa.
+For brevity, PSF variation results are shown in Appendix C.2.
+Each experiment spans a 10-minute detection window, with attacks initiated mid-run.
+For HAI, the attack window includes the actual 164s event and a 3-minute buffer to account for the non-linear model’s delayed anomaly convergence, whereas in ORNL-PS, the attack persists until the end of the run.
+We test PSF values of 8, 10, and 12, and $D$ values of 4, 8, 16, and 32.
 Figures 3 and 4 demonstrate the detection quality with varying D and while fixing the PSF values to 8, 10 and 12 respectively.
@@
 In each case we can observe that the detection consistently occurs within the designated attack duration window providing similar quality as the non-ZKSNARK version.
-In Figures 5 and 6, we depict the detection quality with respect to varying PSF values for distinct values of D.
+Figure 5: Computational error check for SVD of Kalman gain and aggregated covariance matrices
+Figure 6: Computational error check for SVD of Kalman gain and aggregated covariance matrices
+6.2 Computational Correctness of zkSTAR output
+The SVD and Kalman gain decompositions in the zkSTAR framework are computed offline.
+As a result, the zkSTAR framework outputs the reconstruction error, the associated threshold and a compute check flag for the SVD and Kalman gain matrices.
+We define the computational correctness as the absolute error emanating from the in-circuit reconstruction of the SVD and Kalman gain matrices computed with respect to the circuit encoded thresholds.
+Therefore, these experiments demonstrate the variation in relative absolute errors of all matrix elements with respect to the given threshold.
+In our experiments the threshold values for Kalman and SVD checks were set to 0.1.
+Figure 5 shows box plots of relative error for varying D values on the HAI and ORNL-PS datasets (PSF = 12) based on the SVD of predicted covariance matrices.
+For both datasets, relative errors remain below 1 in most cases, indicating reconstruction errors under 0.1.
+Slight error spikes above 1 occur during attack windows due to increased residual volatility, yet even in the worst case, relative errors stay below 1.4 (HAI) and 1.2 (ORNL-PS).
+Figure 6 shows scatter plots of relative errors across varying $D$ values.
+Unlike the SVD check, the Kalman gain check, which is performed every $D$ steps, exhibits an initial staggered trend.
+For the HAI dataset, errors remain consistently low around 1e51e-5, while for the ORNL-PS dataset they are slightly higher yet stable, around 1e21e-2.
+6.3 System Performance of Utility Prover
+We begin by analyzing the system performance of zkSTAR proving system.
+Specifically, we examine the witness, proof, key and circuit sizes in addition to analyzing the utility’s system performance of witness proof generation.
+Dataset PK VK PF
+HAI 16011.879 (2271.245) 3.878 (1.014) 0.039 (0.004)
+ORNL-PS 16516.104 (2735.173) 4.103 (1.221) 0.040 (0.006)
+Table I: TC Circuit: Proof (PF), Proving Key (PK), and Verification Key (VK) sizes for TC (mean(std) in MB)
+Dataset PK VK witness Circuit Size PF
+HAI 12931 3.002 0.023 0.009 0.037
+ORNL-PS 12931 3.002 0.016 0.008 0.035
+Table II: SC Circuit Proof (PF), Proving Key (PK), and Verification Key (VK) sizes for SC in MB)
+D HAI ORNL-PS
+Value Witness Circuit Witness Circuit
+1 0.043(0.001) 0.25 0.036(0.002) 0.24
+4 0.073(0.001) 0.78 0.064(0.001) 0.78
+8 0.114(0.001) 1.50 0.101(0.001) 1.49
+16 0.197(0.001) 2.94 0.176(0.002) 2.93
+32 0.362(0.002) 5.82 0.325(0.004) 5.80
+Table III: Temporal Consistency Witness Size with Varying D (mean(std) in MB)
+Analysis of Proof and Key Sizes: Tables I and II presents the sizes of the proof, proving and verification keys for the temporal consistency and statistical consistency proofs.
+Table II also presents the witness size pertaining to the statistical consistency proofs.
+Additionally, we present the witness and the compiled circuit size for temporal consistency proofs for PSF value 8 separately in Table III.
+In Table I, we can observe that for the case of temporal consistency, the proof size as well as proving and verification key sizes are similar for both datasets.
+On the other hand, in Table II, the circuit, proving and verification key size for the statistical consistency proofs were constant as expected and minor deviations are observed for the proof and witness sizes across both datasets.
+However, we observe an interesting trend with respect to witness and circuit sizes as represented in Table III.
+Mean witness sizes increase gradually from around 43kB to 360kB for HAI and 36kB to 325kB for ORNL-PS.
+On the other hand, the circuit sizes increase from a smaller value of 0.25MB at D value of 1 to 5.8MB for HAI while exhibiting a near identical trend for ORNL-PS dataset as well.
+This trend is natural and expected since increase in D value corresponds to larger TC interval sizes making the TC circuit grow exponentially.
+PSF Hypothesis Test D=1 D=4 D=8 D=16
+Value Time (s) Time (s) Speedup Time (s) Speedup Time (s) Speedup Time (s) Speedup
+8 56.06(22.60) 70.05(3.03) 1.00 80.81(4.19) 3.5 70.61(2.87) 7.93 78.25(3.75) 14.32
+10 54.26(19.17) 69.26(3.22) 1.00 71.32(3.22) 3.89 75.97(8.88) 7.3 77.11(5.01) 14.37
+12 53.49(6.60) 69.89(3.77) 1.00 66.50(2.34) 4.20 74.59(3.19) 7.5 77.79(3.90) 14.37
+Table IV: HAI Dataset: Speedup and Proof Generation Times (mean (std) in seconds)
+PSF Hypothesis Test D=1 D=4 D=8 D=16
+Value Time (s) Time (s) Speedup Time (s) Speedup Time (s) Speedup Time (s) Speedup
+8 49.01 (9.68) 81.74 (2.85) 1.00 57.75 (1.35) 5.66 67.69 (1.74) 9.66 72.64 (1.77) 18.0
+10 50.79 (1.32) 69.82 (2.29) 1.00 65.06 (1.17) 4.29 68.66 (1.75) 8.13 73.28 (1.83) 15.24
+12 52.40 (37.36) 69.27 (3.99) 1.00 64.45 (1.26) 4.29 67.21 (1.22) 8.24 73.32 (2.17) 15.11
+Table V: ORNL-PS Dataset: Speedup and Proof Generation Times (mean (std) in seconds)
+Computational Analysis of Proving System: We analyze the computational performance for temporal and statistical consistency proof generation for both datasets for varying D values.
+Circuit generation and setup costs are presented separately in Appendix C.4.
+Our results are presented in Tables IV and V for HAI and ORNL-PS datasets respectively.
+The data for these experiments are obtained from multiple runs of varying PSF values and represent the mean and standard deviation for one proof invocation of the circuit.
+We also provide a speed up value that represents the average case speedup that can be expected when sequentially generating proofs for the entire horizon.
+We provide the speedup formula and its associated explanation using Equation (48) in Appendix C.1 respectively.
+The temporal consistency proof generation times are represented by the mean and standard deviation values for D values of 1,4,8 and 16.
+The proving times are in general stable with minor variations observed with varying D values.
+We observe that in case of HAI, the mean proof generation time increases slightly going from D values of 1 to 4 before decreasing slightly from 4 to 8 and finally increasing from 8 to 16.
+On the other hand the trend is inverted for ORNL-PS, decreasing slightly from 1 to 4, increasing from 4 to 8 and then decreasing again from 8 to 16.
+Interestingly, there is more volatility for PSF value of 8 in case of either dataset as compared to 10 and 12.
+The non-monotonous trends depicted in Tables IV and V can be due to the non-linear proving costs afforded by ezkl backend which could be explained on the basis of row domain padding used by Halo2 proving systems.
+Further, analyzing speedups, we note a healthy increase between each set of D values regardless of the PSF value or the dataset used.
+Since the circuit size is known to influence the proving times, the trends in Tables IV and V are exactly aligned with those depicted in Table III.
+The speedup value insights point to the fact that having a higher TC interval size (or D value) is almost always more computationally efficient than invoking proof generation on a finer granularity corresponding to a lower TC interval size.
+Lastly, the hypothesis test columns in Tables IV and V reflect a relatively stable values of average proving time corresponding to statistical consistency proofs.
+These values are slightly higher for HAI compared to ORNL-PS which might be explained on the basis of the difference in dimensionality between both problems.
+Figure 7: Witness Generation Performance with varying D for HAI dataset
+Figure 8: Prover Performance for Temporal Consistency with varying D for HAI dataset
+System Performance of Witness and Proof Generation: Figures 7 and 8 present trends pertaining to the CPU utilization and memory consumption for only the HAI dataset.
+We provide the corresponding trends for ORNL-PS in Appendix C.3 due to space constraints.
+The results show a steady and predictable scaling trend, with CPU utilization peaking around 1600%, demonstrating effective parallelization, and memory usage reaching approximately 55 GB, well within available system capacity.
+From Figure 7, we can see greater volatility in CPU utilization and memory consumption for D value of 1 eventually stabilizing for D values of 8 and 16.
+The underlying reason for this trend is that lower values of D demand more frequent proving steps that results in higher volatility in CPU utilization and memory consumption.
+As D grows to 8 and 16, the proving time stabilizes on account of less frequent proving steps encountered during the simulation horizon.
+These results confirm that zkSTAR efficiently leverages multi-core resources with stability in memory consumption under increasing computational load.
+Verification Latency: To gauge the regulatory overhead in terms of verification, we measure the verification times incurred by the regulator for verifying TC and SC proof artifacts.
+In almost all cases, zkSTAR SC and TC proofs are successfully verified in under 1.5s for both datasets with varying TC interval sizes.
+Detailed results pertaining to verification times for each case has been presented in Appendix C.5.
+These results assert the practicality of zkSTAR for near real-time regulatory audits with minimal overhead.
+7 Conclusion
+In this paper, we address the problem of regulatory compliance in data-driven cyberattack detection for utility stakeholders within large-scale Critical Infrastructure Networks (CINs).
+Privacy and efficiency constraints make full-fledged ICS data audits impractical, creating a key challenge for regulators seeking trustworthy verification of local detection outcomes.
+The absence of private, verifiable mechanisms can degrade network-wide situational awareness and increase false positives, resulting in unnecessary downtimes.
+To overcome these limitations, we present zkSTAR, a zkSNARK-based framework that enables publicly verifiable attack detection while preserving complete zero knowledge of local ICS state-space dynamics.
+zkSTAR provides formal guarantees of temporal and statistical consistency for utility-level detections, linking proof verification directly to the authenticity of attack alarms.
+We theoretically prove knowledge-soundness claims in zkSTAR while demonstrating its robustness to attack suppression in case of compromised utilities.
+For scalability, the framework decomposes proof generation into distinct zkSNARK-based kernels, ensuring computational efficiency and modularity.
+We evaluate zkSTAR on a containerized testbed using real-world datasets from HAI and ORNL-PS, demonstrating robust performance across varying TC intervals and precision scales.
+System-level analysis further confirms the framework’s stability in terms of CPU and memory utilization.
+Overall, our results show that zkSTAR enables secure, privacy-preserving, and publicly verifiable cyberattack detection suitable for diverse CIN deployment scenarios.
+Ethics Consideration
+None
+LLM Usage Considerations
+LLMs were used solely for editorial purposes in the preparation of this manuscript (e.g., grammar refinement, stylistic polishing, and clarity improvements).
+All outputs were manually reviewed, verified, and edited by the authors to ensure technical accuracy and originality.
+No new ideas, data, or analyses were introduced through the use of LLMs.
+The literature review, experimental design, results interpretation, and all conceptual contributions were conducted entirely by the authors.
+Appendix A Extended Kalman Filter Model and Training
+A.1 EKF Modeling for ICS
+Our state space modeling framework is based on the extended Kalman filter (EKF) modeling paradigm that is capable of handling non-linearities inherent in the state space process [coskun2017long].
+In order to represent the EKF model, we use a deep neural network driven strategy that consists of one feed-forward and three LSTM sub-modules.
+Specifically, we consider a sensor-driven non-linear system at time $t$, where $x_{t}\in R^{m}$ represents the latent space embedding, $u_{t}\in R^{m}$ represents the control action and $y_{t}\in R^{d}$ represents noisy sensor measurements from asset sensors.
+$x_{t+1}=g(x_{t-1},u_{t-1})+v_{t},$ (23)
+$y_{t}=h(x_{t})+w_{t},$ (24)
+In Equations (23), (24), g,h{g},{h} are the state transition and observation functions respectively.
+The process and measurement noises at time $t$ are denoted by $v_{t}\in R^{m}$, and $w_{t}\in R^{m}$ respectively.
+The process and measurement noises follow multivariate normal distributions with zero mean implying that $v_{t}\sim N(0,Q_{t}),w_{t}\sim N(0,R_{t})$, where $Q_{t},R_{t}$ represent the covariance matrices respectively.
+Such a type of modeling framework has also been used extensively in prior art [smith2011decoupled, mo2014detecting, teixeira2010cyber].
+Figure 9 represents the dynamics of our LSTM based deep neural network that is specifically tailored to establish the extended Kalman filter model for ICS subsystems.
+Figure 9: Non Linear Kalman Filter based Temporal State Estimation
+A.2 EKF training methodology
+During the training phase, the efficacy of the model is assessed by comparing the estimated state $x_{t|t}$ against the ground-truth state $x_{t}$ through a carefully constructed loss function.
+This function typically comprises a mean squared error (MSE) term to minimize the divergence between the estimated and true states.
+To ensure statistical robustness and numerical stability, regularization terms are incorporated to constrain the norms and conditioning of the covariance matrices $Q_{t}$ and $R_{t}$, as well as the Kalman gain $K_{t}$.
+These constraints mitigate overfitting and foster consistent learning behavior across iterations.
+The total loss is subsequently backpropagated through the model’s architecture via automatic differentiation techniques, enabling the optimization of all learnable parameters.
+This includes parameters embedded within the state transition function, the measurement function, and the covariance estimation modules.
+Through iterative training over sequential data, the KFLSTM model progressively refines its ability to encapsulate the system’s dynamic behavior and uncertainty propagation, aligning seamlessly with the theoretical underpinnings of non-linear Kalman filtering.
+Appendix B Theorem and Lemma Proofs
+B.1 Proof of Theorem 4.1
+Proof.
+We know from Definition 3, for practical purposes $\mathcal{H}^{out}_{t}and\mathcal{H}^{in}_{t+1}$ can be used to uniquely identify $x^{out}_{t}$ and $x^{in}_{t+1}$.
+Therefore, it is sufficient to demonstrate the following
+$\mathcal{H}^{in}_{t+1}=\mathcal{H}^{out}_{t}\implies\Phi_{t+1}=1$ (25)
+$\Phi_{t+1}=1\implies\mathcal{H}^{in}_{t+1}=\mathcal{H}^{out}_{t}$ (26)
+[…]
+$\tilde{\Pi}_{t+1}=\texttt{Prove}(pk,\theta,\mathcal{M},\tilde{\Delta}^{in}_{t+1},\tilde{\Delta}^{out}_{t+1})$ (27)
+$\tilde{\Phi}_{t+1}=\texttt{Verify}(vk,\tilde{\Pi}_{t+1},\tilde{\Delta}^{in}_{t+1},\tilde{\Delta}^{out}_{t})$ (28)
+On the other hand, using $\mathcal{H}^{out}_{t}$, for verification purpose we can also construct $\Delta^{in}_{t+1}=(\mathcal{H}^{out}_{t},\tilde{x}^{in}_{t+1},y_{t})$,
+$\Phi_{t+1}=\texttt{Verify}(vk,\tilde{\Pi}_{t+1},\Delta^{in}_{t+1},\tilde{\Delta}^{out}_{t})$ (29)
+However, given $vk_{t}=vk_{t+1}$ and $pk_{t}=pk_{t+1}$, we know that $\Phi_{t+1}\neq\tilde{\Phi}_{t+1}$. As a result $\Phi_{t+1}\neq 1$ which presents a contradiction to our stated assumptions.
+Case 2: Assume that for some $t+1$, $\Phi_{t+1}\neq 1$ such that $\mathcal{H}^{in}_{t+1}=\mathcal{H}^{out}_{t}$.
+Further let $\tilde{\Delta}^{in}_{t+1}=(\mathcal{H}^{out}_{t},x^{in}_{t+1},y_{t+1})$ and $\Delta^{in}_{t+1}=(\mathcal{H}^{in}_{t+1},x^{out}_{t},y_{t+1})$.
+Using Definition 2, we can assert that $\Delta^{in}_{t+1}$ and $\tilde{\Delta}^{in}_{t+1}$ are both valid inputs to $\mathcal{M}$ parametrized by $\theta$.
+Additionally, as a result of the zk-SNARK verification property stated in Definition 6, we can also assert without loss of generality that there exist $\Pi_{t+1},\Delta^{out}_{t+1},\tilde{\Pi}_{t+1},\tilde{\Delta}^{out}_{t+1}$ such that
+$\Phi_{t+1}=\texttt{Verify}(vk,\Pi_{t+1},\Delta^{in}_{t+1},\Delta^{out}_{t+1})=0$ (30)
+$\tilde{\Phi}_{t+1}=\texttt{Verify}(vk,\tilde{\Pi}_{t+1},\tilde{\Delta}^{in}_{t+1},\tilde{\Delta}^{out}_{t+1})=1$ (31)
+Verification outcomes $\Phi_{t+1}=0$ and $\tilde{\Phi}_{t+1}=1$ will only be obtained in case of the following scenarios or a combination thereof.
+Scenario 1 $\Delta^{out}_{t+1}\neq\tilde{\Delta}^{out}_{t+1}$]: We know from Definition 2 that $\Delta^{out}_{t+1},\tilde{\Delta}^{out}_{t+1}$ are deterministic outcomes from $\mathcal{M}$ parametrized by $\theta$.
+Therefore, it can be established that $\Delta^{out}_{t+1}\neq\tilde{\Delta}^{out}_{t+1}$ is a direct consequence of differing inputs to $\mathcal{M}$ parametrized by $\theta$.
+As a result, for this case, it holds that $\tilde{\Delta}^{in}_{t+1}\neq\Delta^{in}_{t+1}$ and as a consequence $\tilde{\Pi}_{t+1}\neq\Pi_{t+1}$ will also hold.
+Scenario 2 $\tilde{\Pi}_{t+1}\neq\Pi_{t+1}$]: We know that the following relationship must hold if $\tilde{\Phi}_{t+1}=1$
+$\tilde{\Pi}_{t+1}=\texttt{Prove}(pk,\theta,\mathcal{M},\tilde{\Delta}^{in}_{t+1},\tilde{\Delta}^{out}_{t+1})$ (32)
+Since, $\tilde{\Pi}_{t+1}\neq\Pi_{t+1}$, we can derive that either $\tilde{\Delta}^{in}_{t+1}\neq\Delta^{in}_{t+1}$ or $\Delta^{out}_{t+1}\neq\tilde{\Delta}^{out}_{t+1}$.
+We have already shown that $\Delta^{out}_{t+1}\neq\tilde{\Delta}^{out}_{t+1}$ directly results from $\tilde{\Delta}^{in}_{t+1}\neq\Delta^{in}_{t+1}$.
+As a result, it conclusively holds that $\tilde{\Delta}^{in}_{t+1}\neq\Delta^{in}_{t+1}$.
+As a result, we establish that divergent verification outcomes $\Phi_{t+1}=0$ and $\tilde{\Phi}_{t+1}=1$ necessarily imply $\tilde{\Delta}^{in}_{t+1}\neq\Delta^{in}_{t+1}$.
+Therefore, $\mathcal{H}^{out}_{t}\neq\tilde{\mathcal{H}}^{in}_{t+1}$ and $x^{in}_{t+1}\neq\tilde{x}^{out}_{t}$ which is a contradiction to our stated assumption.
+∎
+B.2 Proof of Lemma 1
+Proof.
+We can see that under conditions of probabilistic similarity, the following holds
+[…]
+From Section 3.2, we know that this leads to $\tilde{\rho}_{t}=\rho_{t}$ respectively thereby establishing the sufficiency argument.
+[…]
+B.3 Proof of Theorem 4.2
+Proof.
+We know from Theorem 4.1 that $\mathcal{H}^{in}_{t+1}=\mathcal{H}^{out}_{t}\text{ , and }\Phi_{t+1}=1$ is a necessary and sufficient condition for temporal consistency.
+Therefore, to prove this theorem, it is sufficient to show that for a temporally consistent system $\Phi_{t+1}=1\iff\text{Statistical consistency}$.
+(Case 1 $\Phi_{t+1}=1\implies\text{Statistical consistency}$): Assume that $\rho^{real}_{t+1}\neq\rho^{rep}_{t+1}$ but $\Phi_{t+1}=1$.
+Based on Lemma 1, we know that there are two probabilistically dissimilar sensor measurement realizations $y^{real}_{t+1},y^{rep}_{t+1}$ corresponding to $\rho^{real}_{t+1},\rho^{rep}_{t+1}$ respectively.
+This implies that there exists a system of proof artifacts pertaining to $\rho^{rep}_{t+1}$ that satisfy the following relations.
+$\tilde{\Delta}^{in}_{t+1}=(\tilde{\mathcal{H}}^{in}_{t+1},\tilde{x}^{in}_{t+1},y^{rep}_{t+1})$ (34)
+$\tilde{\Delta}^{out}_{t+1}=(\mathcal{H}^{out}_{t+1},\rho^{rep}_{t+1},{x}^{in}_{t+1},r_{t+1},K_{t+1},Q_{t+1},R_{t+1})$ (35)
+$\tilde{\Pi}_{t+1}=\texttt{Prove}(pk,\theta,\mathcal{M},\tilde{\Delta}^{in}_{t+1},\tilde{\Delta}^{out}_{t+1})$ (36)
+$\Phi_{t+1}=\texttt{Verify}(vk,\tilde{\Pi}_{t+1},\tilde{\Delta}^{in}_{t+1},\tilde{\Delta}^{out}_{t+1})=1$ (37)
+On the other hand, we can state that for $\rho^{real}_{t+1}$ there also exists a system of proofs such that
+${\Delta}^{in}_{t+1}=(\mathcal{H}^{in}_{t+1},x^{in}_{t+1},y^{real}_{t+1})$ (38)
+${\Delta}^{out}_{t+1}=(\mathcal{H}^{out}_{t+1},\rho^{real}_{t+1},\hat{x}_{t+1|t+1},r_{t+1},K_{t+1},Q_{t+1},R_{t+1})$ (39)
+${\Pi}_{t+1}=\texttt{Prove}(pk,\theta,\mathcal{M},{\Delta}^{in}_{t+1},{\Delta}^{out}_{t+1})$ (40)
+$\Phi^{\prime}_{t+1}=\texttt{Verify}(vk,{\Pi}_{t},{\Delta}^{in}_{t},{\Delta}^{out}_{t+1})=1$ (41)
+We can clearly see that ${\Delta}^{out}_{t+1}\neq\tilde{\Delta}^{out}_{t+1}$, which implies that ${\Pi}_{t+1}\neq\tilde{\Pi}_{t+1}$.
+However, since both ${\Delta}^{out}_{t+1}$ and $\tilde{\Delta}^{out}_{t+1}$ are valid outputs of the pretrained model, we can further claim that ${\Delta}^{in}_{t+1}\neq\tilde{\Delta}^{in}_{t+1}$.
+Further analyzing ${\Delta}^{in}_{t+1},\tilde{\Delta}^{in}_{t+1}$ we note that $\tilde{y}_{t+1}$ is also probabilistically dissimilar to $y_{t+1}$.
+This means that the only viable possibility is that $(\mathcal{H}^{in}_{t+1},x^{in}_{t+1})\neq(\tilde{\mathcal{H}}^{in}_{t+1},\tilde{x}^{in}_{t+1})$.
+However, this possibility contradicts our temporal consistency constraint thereby proving the intended statement.
+(Case 2 $\text{Statistical consistency}\implies\Phi_{t+1}=1$): Assume that $\Phi_{t+1}\neq 1$ where $\rho^{real}_{t+1}=\rho^{rep}_{t+1}$.
+We have a valid output ${\Delta}^{out}_{t+1}$ for an input ${\Delta}^{in}_{t+1}$ such that.
+${\Delta}^{in}_{t+1}=(\mathcal{H}^{in}_{t+1},x^{in}_{t+1},y^{real}_{t+1})$ (42)
+${\Delta}^{out}_{t+1}=(\mathcal{H}^{out}_{t+1},\rho^{real}_{t+1},{x}^{in}_{t+1},r_{t+1},K_{t+1},Q_{t+1},R_{t+1})$ (43)
+$\Phi_{t+1}=\texttt{Verify}(vk,{\Pi}_{t},{\Delta}^{in}_{t},{\Delta}^{out}_{t+1})=0$ (44)
+Alternately, consider a case for a valid output $\tilde{\Delta}^{out}_{t+1}$ for an input $\tilde{\Delta}^{in}_{t+1}$ such that,
+$\tilde{\Delta}^{in}_{t+1}=(\tilde{\mathcal{H}}^{in}_{t+1},\tilde{x}^{in}_{t+1},y^{rep}_{t+1})$ (45)
+$\tilde{\Delta}^{out}_{t+1}=(\mathcal{H}^{out}_{t+1},\rho^{rep}_{t+1},{x}^{in}_{t+1},r_{t+1},K_{t+1},Q_{t+1},R_{t+1})$ (46)
+$\tilde{\Phi}_{t+1}=\texttt{Verify}(vk,\tilde{\Pi}_{t+1},\tilde{\Delta}^{in}_{t+1},\tilde{\Delta}^{out}_{t+1})=1$ (47)
+Comparing $\tilde{\Phi}_{t+1}$ and $\Phi_{t+1}$, we can conclusively say that $\tilde{\Delta}^{in}_{t+1}\neq{\Delta}^{in}_{t+1}$.
+Based on Lemma 1, we know that the underlying sensor measurement realizations $y^{real}_{t+1},y^{rep}_{t+1}$ are probabilistically similar since statistical consistency is assumed.
+Since, $\tilde{\Delta}^{in}_{t+1}$ is a valid output of $\mathcal{M}$ and $y^{real}_{t+1},y^{rep}_{t+1}$ are probabilistically similar, we can say that the reason for $\tilde{\Phi}_{t+1}\neq{\Phi}_{t+1}$ is purely because of $(\mathcal{H}^{in}_{t+1},x^{in}_{t+1})\neq(\tilde{\mathcal{H}}^{in}_{t+1},\tilde{x}^{in}_{t+1})$, which is a contradiction of our temporal consistency conditions establishing the soundness of the intended statement.
+∎
+B.4 Proof of Theorem 4.3
+Proof.
+We consider $\mathcal{E}^{\mathcal{U}}$ producing proof $\Pi_{t}$ which results in $\Phi_{t}=1$ with an incorrect witness $\mathcal{E}^{\mathcal{A}}(\Delta_{t}^{in}$.
+Under the knowledge-soundness property of Halo2 [halo2_zcash_2023], we can state that
+[…]
+$\quad\wedge\ \big((\Delta_{t}^{in},\Delta_{t}^{out}),\,\mathcal{E}^{\mathcal{A}}(\Delta_{t}^{in},\Delta_{t}^{out})\big)\notin\mathcal{R}_{t}\big]\ \leq\ \text{negl}(\lambda)$
+As a result, the knowledge-soundness property of Halo2 contradicts the assumption of detection suppression.
+Figure 10: HAI Dataset: Detection Quality with varying PSF and fixed D
+Figure 11: ORNL-PS Dataset: Detection Quality with varying PSF and fixed D
+We also know that the following claims hold $\forall t\geq t^{attack}$:
+  1. C1
+    Key invariance ($vk_{t}=vk_{t+1}$ and $pk_{t}=pk_{t+1}$) is necessary for TC based on Theorem 4.1
+  2. C2
+    Hash invariance $\mathcal{H}^{in}_{t+1}=\mathcal{H}^{out}_{t}$ and $\Phi_{t}=1$ is a necessary and sufficient condition for TC based on Theorem 4.1.
+  3. C3
+    TC for successive time steps under key invariance implies SC.
+The infeasibility of detection suppression directly implies a violation of $\Phi_{t}=1$ or of TC-SC relation $\mathcal{R}_{t}$.
+Either violation can be attributed to a violation of either $\textbf{C1},\textbf{C2}$ or C3 applicable $\forall t\geq t^{attack}$.
+Therefore, we can clearly establish that infeasibility of detection suppression is equivalent to violation of either the proof verification or the TC-SC relationship.
+Appendix C Additional Experimental Results
+C.1 Proof Setup and Generation Times
+The speedup relation used in Tables IV, V can be given as:
+[…]
+Dataset DVAL CPU (%) Memory (GB)
+HAI 16 1081.584(604.146) 30.135(9.980)
+ORNL-PS 16 1081.584(604.146) 30.135(9.980)
+Table VI: CPU and Memory Utilization (mean(std)) for Statistical consistency
+Figure 12: Witness Generation Performance with varying D for ORNL-PS dataset
+Figure 13: Prover Performance for Temporal Consistency with varying D for ORNL-PS dataset
+The speedup computation is based on obtaining the ratio between the average time consumed for the entire set of D unrolled loop iterations when $D=1$ and the corresponding values when $D\neq 1$.
+Since the time horizon appears in the numerator and denominator, they cancel each other resulting in Equation (48).
+C.2 Detection Quality with varying PSF
+In Figures 10 and 11, we depict the detection quality with respect to varying PSF values for distinct values of D.
 In this case we can observe that the lower PSF values usually exhibit greater volatility in detection, while higher PSF values correspond to a less volatile nature of the test statistic.
@@
 Regardless, the experiments establish the robustness in detection quality of the zkSTAR framework.
-The setup times for HAI and ORNL-PS dataset has been provided in Tables III and IV respectively.
-Figure 7: Computational error check for SVD of Kalman gain and aggregated covariance matrices
-Figure 8: Computational error check for SVD of Kalman gain and aggregated covariance matrices
-Dataset PK VK PF
-HAI 16011.879 (2271.245) 3.878 (1.014) 0.039 (0.004)
-ORNL-PS 16516.104 (2735.173) 4.103 (1.221) 0.040 (0.006)
-Table I: Proof (PF), Proving Key (PK), and Verification Key (VK) sizes for Temporal Consistency (mean(std) in MB)
-Dataset PK VK witness Circuit Size PF
-HAI 12931 3.002 0.023 0.009 0.037
-ORNL-PS 12931 3.002 0.016 0.008 0.035
-Table II: Proof (PF), Proving Key (PK), and Verification Key (VK) sizes for Statistical consistency (mean(std) in MB)
+C.3 Analysis of System Performance of ORNL-PS
+Figures 12 and 13 present the system performance characterized by CPU utilization and memory consumption for ORNL-PS dataset for witness generation and proof generation mechanism.
+In Table VI, we provide the CPU utilization and memory consumption for the statistical consistency proofs for HAI and ORNL-PS datasets.
+C.4 Analyses of Circuit Setup Costs
 Setup Setup Time (s)
 Phase D=1 D=4 D=8 D=16
-Model Export 24.77 3.25 7.96 24.21
-Settings Generation 30.02 3.53 9.55 28.99
-Circuit Compilation 24.46 2.24 7.06 24.90
-Circuit Setup 171.63 146.84 162.04 167.55
-Table III: HAI Setup Time Breakdown (in seconds)
+Model Export 1.07 3.29 7.92 23.88
+Settings Generation 0.74 3.73 9.80 29.20
+Circuit Compilation 0.31 2.31 7.12 24.20
+Circuit Setup 155.04 156.06 174.11 176.53
+Table VII: HAI Setup Time Breakdown (in seconds)
 Setup Setup Time (s)
 Phase D=1 D=4 D=8 D=16
-Model Export 24.77 3.25 7.96 24.21
-Settings Generation 30.02 3.53 9.55 28.99
-Circuit Compilation 24.46 2.24 7.06 24.90
-Circuit Setup 171.63 146.84 162.04 167.55
-Table IV: ORNL-PS Setup Time Breakdown (in seconds)
-D HAI ORNL-PS
-Value Witness Circuit Witness Circuit
-1 0.197(0.001) 2.94 0.176(0.002) 2.93
-4 0.073(0.001) 0.78 0.064(0.001) 0.78
-8 0.114(0.001) 1.50 0.101(0.001) 1.49
-16 0.197(0.001) 2.94 0.176(0.002) 2.93
-32 0.362(0.002) 5.82 0.325(0.004) 5.80
-Table V: Temporal Consistency Witness Size with Varying D (mean(std) in MB)
-PSF Hypothesis Test D=1 D=4 D=8 D=16
-Value Time (s) Time (s) Speedup Time (s) Speedup Time (s) Speedup Time (s) Speedup
-8 56.06(22.60) 90.83(5.41) 1.00 80.81(4.19) 4.50 70.61(2.87) 10.29 78.25(3.75 18.57
-10 54.26(19.17) 76.16(1.55) 1.00 71.32(3.22) 4.27 75.97(8.88) 8.02 77.11(5.01) 15.80
-12 53.49(6.60) 76.74(3.05) 1.00 66.50(2.34) 4.62 74.59(3.19) 8.23 77.79(3.90) 15.78
-Table VI: HAI Dataset: Speedup and Proof Generation Times (mean (std) in seconds)
-PSF Hypothesis Test D=1 D=4 D=8 D=16
-Value Time (s) Time (s) Speedup Time (s) Speedup Time (s) Speedup Time (s) Speedup
-8 49.01 (9.68) 73.40 (1.92) 1.00 57.75 (1.35) 5.08 67.69 (1.74) 8.67 72.64 (1.77) 16.17
-10 50.79 (1.32) 74.00 (2.06) 1.00 65.06 (1.17) 4.55 68.66 (1.75) 8.62 73.28 (1.83) 16.16
-12 52.40 (37.36) 99.94 (1.87) 1.00 64.45 (1.26) 6.20 67.21 (1.22) 11.90 73.32 (2.17) 21.81
-Table VII: ORNL-PS Dataset: Speedup and Proof Generation Times (mean (std) in seconds)
-VI-B Computational Correctness of zkSNARK outputs
-The SVD and Kalman gain decompositions in the zkSTAR framework are computed offline.
-As a result, the zkSTAR framework outputs the reconstruction error, the associated threshold and a compute check flag for the SVD and Kalman gain matrices.
-We define the computational correctness as the absolute error emanating from the in-circuit reconstruction of the SVD and Kalman gain matrices computed with respect to the circuit encoded thresholds.
-Therefore, these experiments demonstrate the variation in relative absolute errors of all matrix elements with respect to the given threshold.
-In our experiments the threshold values for Kalman and SVD checks were set to 0.1.
-Figure 7 denotes box plot series of relative error for varying D values for HAI and ORNL-PS datasets with PSF=12 pertaining to the SVD of the predicted covariance matrices.
-We note that in case of both datasets, the relative error remains well below 1 for a majority of the cases.
-This indicates that the reconstruction error remains less than 0.1 in most cases.
-We observe higher incidence of relative error values going above 1 during the attack window which can be due to the greater volatility in residual values computed during an attack.
-However, even in the worst case, the relative error with respect to matrix elements does not exceed 1.4 and 1.2 respectively for HAI and ORNL datasets.
-Figure 8 denotes scatter plot series of relative errors for varying D values.
-We note that unlike the SVD check, the Kalman gain check is carried out after each D time steps, resulting in a staggered trend for the first few iterations.
-The Kalman gain check for both datasets conclusively demonstrates a consistently stable and low error value in the neighborhood of 1e51e-5 for HAI dataset.
-However, for the ORNL-PS dataset, we see a higher but highly stable relative error values aroun 1e21e-2.
-VI-C System Performance of Utility Prover
-We begin by analyzing the system performance of zkSTAR proving system.
-Specifically, we examine the witness, proof, key and circuit sizes in addition to analyzing the utility’s system performance of witness proof generation.
-Analysis of Proof and Key Sizes: Tables I and II presents the sizes of the proof, proving and verification keys for the temporal consistency and statistical consistency proofs.
-Table II also presents the witness size pertaining to the statistical consistency proofs.
-Additionally, we present the witness and the compiled circuit size for temporal consistency proofs separately in Table V.
-In Table I, we can observe that for the case of temporal consistency, the proof size as well as proving and verification key sizes are similar for both datasets.
-On the other hand, in Table II, the circuit, proving and verification key size for the statistical consistency proofs were constant as expected and minor deviations are observed for the proof and witness sizes across both datasets.
-However, we observe an interesting trend with respect to witness and circuit sizes as represented in Table V.
-It can be seen that the circuit and witness sizes for both datasets exhibit a drop from D values of 1 to 4 before gradually increasing for more extreme D values such as 8, 16 and 32.
-This behavior can be potentially explained on the basis of row domain padding used by Halo2 proving systems [44] (such as ezkl) which pads rows to the nearest value of $2^{k}$.
-This behavior is also evidenced by the fact that the witness and circuit sizes almost double from 4 to 12.
-As a consequence, we can conclude that D=4 might correspond to a sweet spot for the non linear Kalman Filter LSTM based circuit leading to a lower circuit and witness size.
-Analyses of Circuit Setup Costs: In Tables III and IV, we present the time incurred during the setup phase of zkSTAR pertaining to model export, settings generation, circuit compilation and circuit setup for HAI and ORNL-PS datasets.
-In each of these tables we see an almost identical trend as those seen in Table V for each category of setup steps.
-These trends can be explained on the basis of the fact that the setup of zk-SNARKs relies significantly on the size of the circuits themselves.
-Therefore, the trends in Tables III and IV can be a natural consequence of the row domain padding adopted by Halo 2 proving backends.
-This ultimately results in trends wherein a decrease in time consumption is observed from D value of 1 to 4 followed by an increase for more extreme values like 8 and 16.
-Computational Analysis of Proving System: We analyze the computational performance for temporal and statistical consistency proof generation for both datasets for varying D values.
-Our results are presented in Tables VI and VII for HAI and ORNL-PS datasets respectively.
-The data for these experiments are obtained from multiple runs of varying PSF values and represent the mean and standard deviation for one proof invocation of the circuit.
-We also provide a speed up value that represents the average case speedup that can be expected when sequentially generating proofs for the entire horizon.
-We provide the speedup formula and its associated explanation in Equation (58) and Appendix A-A respectively.
-The temporal consistency proof generation times are represented by the mean and standard deviation values for D values of 1,4,8 and 16.
-The proving times are in general stable with minor variations observed with varying D values.
-We observe that the mean proof generation time decreases slightly from D values of 1 to 8, before increasing slightly at 16.
-This trend is consistently observed across both datasets for all PSF values.
-The non-monotonous trends depicted in Tables VI and VII can be due to the non-linear proving costs afforded by ezkl backend.
-Fundamentally, this trend can be explained on the basis of Table V which highlights the non monotonous behavior of witness and circuit sizes resulting from row domain padding used by Halo2 proving systems.
-Since the proving times are directly proportional to the circuit size, the trends in Tables VI and VII are exactly aligned with those depicted in Table V.
-Lastly, the hypothesis test columns in Tables VI and VII reflect a relatively stable values of average proving time corresponding to statistical consistency proofs.
-These values are slightly higher for HAI compared to ORNL-PS which might be explained on the basis of the difference in dimensionality between both problems.
-Figure 9: Witness Generation Performance with varying D for HAI dataset
-Figure 10: Prover Performance for Temporal Consistency with varying D for HAI dataset
-System Performance of Witness and Proof Generation: Figures 9 and 10 present trends pertaining to the CPU utilization and memory consumption for only the HAI dataset.
-we provide the corresponding trends for ORNL-PS in Appendix A-B due to space constraints as well as similar nature of both curves.
-From Figure 9, we can see greater volatility in CPU utilization and memory consumption for D value of 1.
-We notice that this trend gradually stabilizes for D values of 8 and 16.
-The underlying reason for this trend is that proof generation is an inherently lower values of D demand more frequent proving steps that results in higher volatility in CPU utilization and memory consumption.
-As D grows to 8 and 16, the proving time stabilizes on account of less frequent proving steps encountered during the simulation horizon.
-Additionally, we present the proving time system performance using CPU utilization and memory consumption for a random proving window for varying D values.
-We observe a steady uptick in CPU utilization and memory consumption reaching a value of 1600% and 55GB respectively.
-The CPU utilization trends reflect the multi-processing capabilities on the hardware that was used which comprised of 16 vCPUs.
-On the other hand, the memory consumption of 55GB indicates that the proving process consumed more than half the total memory available on the device.
-VII Conclusion
-In this paper, we address the problem of regulatory compliance in data-driven cyberattack detection for utility stakeholders within large-scale Critical Infrastructure Networks (CINs).
-Privacy and efficiency constraints make full-fledged ICS data audits impractical, creating a key challenge for regulators seeking trustworthy verification of local detection outcomes.
-The absence of private, verifiable mechanisms can degrade network-wide situational awareness and increase false positives, resulting in unnecessary downtimes.
-To overcome these limitations, we propose zkSTAR, a zkSNARK-based framework that enables publicly verifiable attack detection while preserving complete zero knowledge of local ICS state-space dynamics.
-zkSTAR provides formal guarantees of temporal and statistical consistency for utility-level detections, linking proof verification directly to the authenticity of attack alarms.
-For scalability, the framework decomposes proof generation into distinct zkSNARK-based kernels, ensuring computational efficiency and modularity.
-We evaluate zkSTAR on a containerized testbed using real-world datasets from HAI and ORNL-PS, demonstrating robust performance across varying temporal consistency intervals and precision scales.
-System-level analysis further confirms the framework’s stability in terms of CPU and memory utilization.
-Overall, our results show that zkSTAR enables secure, privacy-preserving, and publicly verifiable cyberattack detection suitable for diverse CIN deployment scenarios.
-Appendix A Additional Experimental Results
-A-A Proof Setup and Generation Times
-The speedup relation used in Tables VI and VII can be given as:
-[…]
-Dataset DVAL CPU (%) Memory (GB)
-HAI 16 1081.584(604.146) 30.135(9.980)
-ORNL-PS 16 1081.584(604.146) 30.135(9.980)
-Table VIII: CPU and Memory Utilization (mean(std)) for Statistical consistency
-Figure 12: Witness Generation Performance with varying D for ORNL-PS dataset
-Figure 13: Prover Performance for Temporal Consistency with varying D for ORNL-PS dataset
-The speedup computation is based on obtaining the ratio between the average time consumed for the entire set of D unrolled loop iterations when $D=1$ and the corresponding values when $D\neq 1$.
-Since the time horizon appears in the numerator and denominator, they cancel each other resulting in Equation (58).
-A-B Analysis of System Performance of ORNL-PS Dataset
-Figures 12 and 13 present the system performance characterized by CPU utilization and memory consumption for ORNL-PS dataset for witness generation and proof generation mechanism.
-In Table VIII, we provide the CPU utilization and memory consumption for the statistical consistency proofs for HAI and ORNL-PS datasets.
+Model Export 1.51 3.25 7.96 24.21
+Settings Generation 0.74 3.53 9.55 28.99
+Circuit Compilation 0.34 2.24 7.06 24.90
+Circuit Setup 162.31 146.84 162.04 167.55
+Table VIII: ORNL-PS Setup Time Breakdown (in seconds)
+The mean setup times for HAI and ORNL-PS dataset has been provided in Tables VII and VIII respectively for PSF values of 8, 10 and 12.
+These tables report zkSTAR times for model export, settings generation, circuit compilation, and circuit setup on the HAI and ORNL-PS datasets.
+The results exhibit trends similar to Table III, primarily reflecting circuit-size dependencies inherent to zk-SNARK setup.
+C.5 Verification Times
+Table IX: Verification times for HAI and ORNL-PS datasets (PSF=10).
+HAI ORNL-PS
+D Type Time (s) D Type Time (s)
+1 prediction 1.22 1 prediction 1.26
+8 prediction 1.34 8 prediction 1.40
+16 prediction 1.09 16 prediction 1.31
+16 hypothesis 1.24 16 hypothesis 1.46
+Table IX summarizes representative verification latencies for zkSTAR proofs under varying TC interval sizes of 1,8 and 16 for prediction and hypothesis test components.
+Across all scenarios—including both normal and attack conditions—verification completed in approximately 1.0–1.5 s on commodity hardware, with all proofs passing successfully.
+These results confirm that zkSTAR’s zero-knowledge verification incurs minimal overhead and is practical for near–real-time regulatory audits.

### v3

@@
 Such attacks often target industrial control systems (ICSs), degrading asset performance and triggering cascading failures across critical infrastructure networks (CINs).
-While detection occurs locally at utilities, regulatory bodies such as Information Sharing and Analysis Centers (ISACs) aggregate alarms in real time to enhance situational awareness and reduce false positives [dhs_oig2].
+While detection occurs locally at utilities, regulatory bodies such as Information Sharing and Analysis Centers (ISACs) aggregate alarms in real time to enhance situational awareness and reduce false positives [7].
 The effectiveness of these regulators, however, hinges on the integrity and reliability of local detection data.
-Recent real-world incidents [langner2011stuxnet, jeffries2022cyber, salazar2024tale] have demonstrated that adversaries can compromise both IT and OT layers to obscure, delay, or disable detection of underlying anomalies.
+Recent real-world incidents [21, 18, 37] have demonstrated that adversaries can compromise both IT and OT layers to obscure, delay, or disable detection of underlying anomalies.
 Such incidents expose a fundamental regulatory blind spot pertaining to verifying the integrity of reported detections without re-auditing raw operational data.
-Moreover, inconsistent or low-quality local insights can undermine the ISAC’s ability to construct a reliable network-wide situational awareness model, potentially causing unnecessary service disruptions [dhs_oig2].
+Moreover, inconsistent or low-quality local insights can undermine the ISAC’s ability to construct a reliable network-wide situational awareness model, potentially causing unnecessary service disruptions [7].
 To address this gap, we introduce zkSTAR, a framework enabling publicly verifiable detection of data-driven ICS attacks in complete zero knowledge of local ICS state-space dynamics.
@@
 Leveraging knowledge-soundness and provably consistent outcomes, it ensures that compromised utilities cannot falsify or suppress alarms even if their IT–OT stacks are breached.
-The framework enables verifiable regulatory oversight while preserving utility data confidentiality [dhs_oig2], eliminating the need for costly, privacy-intrusive audits that have long impeded secure information sharing [nolan2015cybersecurity].
+The framework enables verifiable regulatory oversight while preserving utility data confidentiality [7], eliminating the need for costly, privacy-intrusive audits that have long impeded secure information sharing [31].
 Additionally, by adopting an on-demand, lazy proof generation paradigm, zkSTAR ensures a low utility overhead as well.
-Conventional ICS anomaly detection relies on state-space models that estimate expected sensor behavior and compute residuals to flag deviations from steady-state conditions [li2020detection].
-Statistical hypothesis tests on these residuals distinguish normal from abnormal operation, triggering alarms when observed measurements significantly diverge from model predictions [li2020deep].
+Conventional ICS anomaly detection relies on state-space models that estimate expected sensor behavior and compute residuals to flag deviations from steady-state conditions [23].
+Statistical hypothesis tests on these residuals distinguish normal from abnormal operation, triggering alarms when observed measurements significantly diverge from model predictions [25].
 With existing approaches, regulators can verify such detection outcomes only through full data audits of ICS operations—requiring utilities to relinquish proprietary IoT datasets and raising severe privacy and data-governance concerns.
@@
 2 Related Works
-Attacks on industrial control systems (ICSs) that target IT layers can often be detected through network traffic monitoring [caselli2016specification, ye2004robustness].
-However, data-driven attacks that manipulate sensor or control data can evade such approaches by simultaneously impacting both IT and OT layers [urbina2016limiting, kang2014cyber, drias2015taxonomy], potentially causing cascading failures across critical infrastructure.
-More recently, there have been several attack vulnerabilities that have been discovered such as like Industroyer and Industroyer2 [salazar2024tale], Triton [jeffries2022cyber] and Stuxnet [langner2011stuxnet].
+Attacks on industrial control systems (ICSs) that target IT layers can often be detected through network traffic monitoring [4, 45].
+However, data-driven attacks that manipulate sensor or control data can evade such approaches by simultaneously impacting both IT and OT layers [42, 19, 8], potentially causing cascading failures across critical infrastructure.
+More recently, there have been several attack vulnerabilities that have been discovered such as like Industroyer and Industroyer2 [37], Triton [18] and Stuxnet [21].
 These vulnerabilities have demonstrated the feasibility of joint manipulation of control systems and monitoring data which have motivated the need for robust verifiable detection mechanisms for ICS.
-To address some of these challenges, several model-based mechanisms have been proposed [huang2018online, hoehn2016detection, rahman2017multi].
-Among these, Kalman filter based approaches are particularly robust and flexible [li2020deep, li2020detection, li2021degradation, ramanan2021blockchain, li2022online], leveraging residuals from state-space models as statistical evidence for anomaly detection.
-These residual-based tests differentiate routine faults from malicious attacks [li2021degradation] and can support decentralized, network-wide alarm dissemination [ramanan2021blockchain].
-From the compliance perspective, however, it is essential for the regulator to receive timely notifications regarding attack detection at utility stakeholders in addition to verifying the integrity of attack detection so as to limit its network-wide impacts and reduce false alarm rates [ramanan2021blockchain, dhs_oig2].
+To address some of these challenges, several model-based mechanisms have been proposed [16, 13, 35].
+Among these, Kalman filter based approaches are particularly robust and flexible [25, 23, 24, 36, 22], leveraging residuals from state-space models as statistical evidence for anomaly detection.
+These residual-based tests differentiate routine faults from malicious attacks [24] and can support decentralized, network-wide alarm dissemination [36].
+From the compliance perspective, however, it is essential for the regulator to receive timely notifications regarding attack detection at utility stakeholders in addition to verifying the integrity of attack detection so as to limit its network-wide impacts and reduce false alarm rates [36, 7].
 Therefore, regulatory compliance goals would ideally enable stakeholders to provide publicly verifiable detection outcomes that are accurate and trustworthy without exposing private operational data.
-Achieving such verifiable compliance under privacy constraints motivates the use of zero-knowledge proofs [narayan2015verifiable], enabling a prover to convince a verifier of correctness without revealing sensitive information.
+Achieving such verifiable compliance under privacy constraints motivates the use of zero-knowledge proofs [29], enabling a prover to convince a verifier of correctness without revealing sensitive information.
 Originating from foundational work by Goldwasser et al.
-[goldwasser2019knowledge] and Blum et al.
-[blum2019non], zero-knowledge proofs evolved into succinct, non-interactive forms (zk-SNARKs) that support efficient verification [kilian1992note, micali2000computationally, gennaro2013quadratic, parno2016pinocchio, ben2013snarks, costello2015geppetto, ben2014succinct].
-Recent applications in machine learning verify training integrity and model ownership [huang2022zkmlaas, wang2022ezdps, nguyen2022preserving, yang2023fedzkp].
+[12] and Blum et al.
+[3], zero-knowledge proofs evolved into succinct, non-interactive forms (zk-SNARKs) that support efficient verification [20, 27, 11, 34, 1, 6, 2].
+Recent applications in machine learning verify training integrity and model ownership [14, 43, 30, 44].
 While existing zk-SNARK frameworks enable general-purpose verifiable computation, none target the verification of dynamical system integrity or hypothesis-test correctness which are cornerstones of industrial attack detection.
-Recent works on verifiable differential privacy [narayan2015verifiable] and secure multiparty analytics [froelicher2020drynx] focus primarily on data confidentiality but not on verifiable detection correctness, a gap zkSTAR directly addresses.
+Recent works on verifiable differential privacy [29] and secure multiparty analytics [10] focus primarily on data confidentiality but not on verifiable detection correctness, a gap zkSTAR directly addresses.
 To alleviate this core critical gap, this paper introduces zkSTAR, a framework that enables provable, privacy-preserving verification of Kalman filter–based detection mechanisms, paving the way for regulatory compliance in CINs.
 3 Nonlinear State Space Modeling
-For developing zkSTAR, we leverage the extended Kalman Filter (EKF) state space models owing to their flexibility in handling highly non-linear spatial and temporal interdependencies among several ICS components [coskun2017long].
+For developing zkSTAR, we leverage the extended Kalman Filter (EKF) state space models owing to their flexibility in handling highly non-linear spatial and temporal interdependencies among several ICS components [5].
 Therefore, we characterize the EKF-based, utility-level attack detection model as a foundational component of zkSTAR for achieving regulatory compliance under zero knowledge.
@@
 3.2 Hypothesis Test for Anomaly Detection
-We consider residual covariance $S_{t}=U_{t}\Sigma_{t}U_{t}^{T}$ coupled with residual measurements $r_{t}$ for detecting covert data-driven attacks [li2020detection].
+We consider residual covariance $S_{t}=U_{t}\Sigma_{t}U_{t}^{T}$ coupled with residual measurements $r_{t}$ for detecting covert data-driven attacks [23].
 Without loss of generality we note that it is also possible to apply our techniques for detecting false data injection by monitoring $R_{t}$.
-We denote the standardized vector of PC scores as $\tau_{t}=(S_{t})^{-1/2}(r_{t})=U_{t}\Sigma_{t}^{-1/2}(r_{t})$ such that $\tau_{t}\sim N(0,I)$ [li2020detection, hwang2017chi].
-This implies that $\tau_{t}||^{2}_{2}\sim\chi^{2}_{p}$, where $p<d$ is the degrees of freedom corresponding to the number of principal components used [hwang2017chi].
+We denote the standardized vector of PC scores as $\tau_{t}=(S_{t})^{-1/2}(r_{t})=U_{t}\Sigma_{t}^{-1/2}(r_{t})$ such that $\tau_{t}\sim N(0,I)$ [23, 17].
+This implies that $\tau_{t}||^{2}_{2}\sim\chi^{2}_{p}$, where $p<d$ is the degrees of freedom corresponding to the number of principal components used [17].
 $\rho=\begin{cases}1,&\text{if }T_{\chi^{2},t}=\|\tau_{t}\|^{2}_{2}>\chi^{2}_{m,\alpha},\\ 0,&\text{otherwise.}\end{cases}$ (8)
-A generic hypothesis testing procedure applicable to sensor and residual monitoring has been discussed in greater detail in [li2021degradation].
+A generic hypothesis testing procedure applicable to sensor and residual monitoring has been discussed in greater detail in [24].
 We define a measure of probabilistic similarity to help establish the distinctness of two sensor measurement vectors as stated in Definition 1.
@@
 Definition 2 portrays a hashing function that will be utilized to ensure the integrity of state space estimates.
-We assume that the function hash preserves integrity, one-wayness, hiding and collision resistance properties [menezes1996handbook].
+We assume that the function hash preserves integrity, one-wayness, hiding and collision resistance properties [26].
 Definition 3 (Pretrained State Space Model).
@@
 Moreover, with the help of Lemma 1, we can also state that a data driven ICS attack orchestration might in theory go undetected under the zkSTAR framework when the malicious sensor measurements are probabilistically similar to the originally recorded sensor measurements.
-Practically, such attacks end up being inconsequential [li2021degradation] since they are unable to effect any material damage to the ICS assets.
+Practically, such attacks end up being inconsequential [24] since they are unable to effect any material damage to the ICS assets.
 Therefore, Lemma 1 serves as the basis for Theorem 4.2 which derives an integral link between key invariance, statistical and temporally consistency.
@@
 Specifically, we design zkSNARK based kernels that target individual sub tasks pertaining to TC and SC attributes of EKF predictions and detection outcomes capable of operating on batched EKF residual outputs.
-We note that current zkSNARK frameworks such as ezkl lack support for linear solvers, SVD decompositions, and derivative evaluations of complex sub modules due to their high computational cost in zero knowledge [ezkl2024].
+We note that current zkSNARK frameworks such as ezkl lack support for linear solvers, SVD decompositions, and derivative evaluations of complex sub modules due to their high computational cost in zero knowledge [40].
 To address these limitations, we employ a compute-offline–validate-online strategy, performing such expensive operations off-circuit while verifying their consistency with the state-space model within the zkSNARK circuit.
@@
 The zkSTAR framework relies on several assumptions pertaining to the properties of the zero-knowledge proving system.
-First, since zkSTAR uses the Halo2 proof generation paradigm as a consequence of using the ezkl framework [ezkl2024], it implicitly depends on elliptic-curve cryptographic primitives and on the integrity of the Common Reference String (CRS).
+First, since zkSTAR uses the Halo2 proof generation paradigm as a consequence of using the ezkl framework [40], it implicitly depends on elliptic-curve cryptographic primitives and on the integrity of the Common Reference String (CRS).
 As a result, we assume that the CRS is generated through a publicly verifiable ceremony per-utility.
-Second, we assume that the EKF model is well-formed and is strongly sensitive to the Jacobian matrices [huang2008analysis] implying that slight deviations in SVD or Kalman gain computations would lead to inconsistent state and residual estimates violating either TC or SC conditions.
-Third, we assume that verifier is honest-but-curious and that public values of the proof artifacts do not violate privacy of utility state space which is a common requirement of existing regulators [doe].
+Second, we assume that the EKF model is well-formed and is strongly sensitive to the Jacobian matrices [15] implying that slight deviations in SVD or Kalman gain computations would lead to inconsistent state and residual estimates violating either TC or SC conditions.
+Third, we assume that verifier is honest-but-curious and that public values of the proof artifacts do not violate privacy of utility state space which is a common requirement of existing regulators [9].
 Finally, we assume that communication channels are assumed to preserve integrity but not confidentiality which can be alleviated by using transport-level encryption like TLS for privacy during proof transmission.
@@
 Figure 4: ORNL-PS Dataset: Detection Quality with varying D and fixed PSF
-For experimental studies, we utilize the HAI [hai] and ORNL-PS [pan2015classification, pan2015developing] datasets.
+For experimental studies, we utilize the HAI [38] and ORNL-PS [32, 33] datasets.
 To build an asset level anomaly detector, we generate pretrained non-linear models using the EKF approach to accurately reflect the state-space of the critical asset subsystems as described in Section 3.
@@
 System Implementation Details: All experiments were carried out on a virtual machine (VM) running Ubuntu 24.04 with 100GB of RAM and 16 vCPUs using Python 3.11 with the detection model inference using PyTorch 2.7.1.
-We utilize the ezkl library [ezkl2024] for generating all the zkSNARKs associated with the zkSTAR framework.
+We utilize the ezkl library [40] for generating all the zkSNARKs associated with the zkSTAR framework.
 We implement zkSTAR using both native and containerized execution modes.
@@
 Overall, our results show that zkSTAR enables secure, privacy-preserving, and publicly verifiable cyberattack detection suitable for diverse CIN deployment scenarios.
-Ethics Consideration
-None
-LLM Usage Considerations
-LLMs were used solely for editorial purposes in the preparation of this manuscript (e.g., grammar refinement, stylistic polishing, and clarity improvements).
-All outputs were manually reviewed, verified, and edited by the authors to ensure technical accuracy and originality.
-No new ideas, data, or analyses were introduced through the use of LLMs.
-The literature review, experimental design, results interpretation, and all conceptual contributions were conducted entirely by the authors.
 Appendix A Extended Kalman Filter Model and Training
 A.1 EKF Modeling for ICS
-Our state space modeling framework is based on the extended Kalman filter (EKF) modeling paradigm that is capable of handling non-linearities inherent in the state space process [coskun2017long].
+Our state space modeling framework is based on the extended Kalman filter (EKF) modeling paradigm that is capable of handling non-linearities inherent in the state space process [5].
 In order to represent the EKF model, we use a deep neural network driven strategy that consists of one feed-forward and three LSTM sub-modules.
@@
 The process and measurement noises follow multivariate normal distributions with zero mean implying that $v_{t}\sim N(0,Q_{t}),w_{t}\sim N(0,R_{t})$, where $Q_{t},R_{t}$ represent the covariance matrices respectively.
-Such a type of modeling framework has also been used extensively in prior art [smith2011decoupled, mo2014detecting, teixeira2010cyber].
+Such a type of modeling framework has also been used extensively in prior art [39, 28, 41].
 Figure 9 represents the dynamics of our LSTM based deep neural network that is specifically tailored to establish the extended Kalman filter model for ICS subsystems.
@@
 Proof.
-We consider $\mathcal{E}^{\mathcal{U}}$ producing proof $\Pi_{t}$ which results in $\Phi_{t}=1$ with an incorrect witness $\mathcal{E}^{\mathcal{A}}(\Delta_{t}^{in}$.
-Under the knowledge-soundness property of Halo2 [halo2_zcash_2023], we can state that
+For detection suppression, we assume that the compromised utility possesses an extractor $\mathcal{E}^{\mathcal{U}}$ that produces a valid proof $\Pi_{t}$ which results in $\Phi_{t}=1$ with an incorrect witness $\mathcal{E}^{\mathcal{A}}(\Delta_{t}^{in},\Delta_{t}^{out})\neq\Omega_{t}$, with a non-negligible probability.
+We consider $\mathcal{E}^{\mathcal{U}}$ producing proof $\Pi_{t}$ which results in $\Phi_{t}=1$ with an incorrect witness $\mathcal{E}^{\mathcal{A}}(\Delta_{t}^{in},\Delta_{t}^{out})$.
+Under the knowledge-soundness property of Halo2 [46], we can say
 […]